### v2

# Quantum Mechanics Relative to a Quantum Reference System: a Relative State Approach

Abstract
This paper proposes an intrinsic or background-independent quantum framework based on entangled state rather than absolute quantum state, it describes a quantum relative state between the under-study quantum system and the quantum measuring apparatus as a quantum reference system, without relying on any external absolute parameter. The paper focuses on a simple example, in which a quantum object’s one-dimensional position as an under-study quantum system, and a quantum clock as a quantum reference system or quantum measuring apparatus. The evolution equation of the state of the quantum object’s position with respect to the state of the quantum clock is given coming from the Ricci-flat Kähler-Einstein equation. In a linear and non-relativistic approximation, the framework recovers the equation of the standard quantum mechanics, in which an intrinsic potential related to some “inertial force” is automatically incorporated in the covariant derivative. A physical relative probability interpretation and a geometric non-trivial fiber bundle interpretation of the entangled state in this intrinsic quantum framework are given. Furthermore, some non-inertial effects, such as the “inertial force”, coming from the general covariance of the intrinsic quantum framework are also discussed. Compared with the functional integral approach which is more easily to generalize the quantum clock to the quantum spacetime reference frame and study quantum gravity, the relative state approach as a canonical description is more suitable for conceptually demonstrating the connections to the standard formalism and interpretation of the quantum mechanics.
I Introduction
The formulation and interpretation of quantum theory in its current form are based on the concept of classical inertial frames. The Schrödinger equation governing the evolution of quantum states in its present form is valid only in classical inertial frames that employ Newtonian time as a global absolute parameter. One of the key motivations behind the development of General Relativity was the absence of the global absolute time due to the constancy of the speed of light, which necessitated the removal of the special status of the inertial frames in physics and the establishment of physical theories (of cause should includes quantum theory) on concepts that hold true in general coordinate systems. In geometry, the evolution began from the analytical geometric approach using a global Cartesian coordinate system, to Gauss’s realization that measuring from local coordinate systems on two-dimensional surface suffices to define its intrinsic geometry, without the need to embed it into a three-dimensional space. This was further generalized by Riemann, who extended Gauss’s ideas on intrinsic geometry through the use of a local quadratic distance form. Starting with Gauss and culminating in Riemann, the method of intrinsic geometry, independent of the coordinate systems, was pioneered. This coordinate-independent spirit or program reached its physical pinnacle in Einstein’s classical General Relativity, bringing about a historic synthesis in classical physics. However, this spirit of the intrinsic geometry has not been fully incorporated into quantum theory. Relative to the principles of the intrinsic geometry and General Relativity, the current framework of quantum theory still requires and relies on the description in terms of an external absolute inertial coordinate system. It remains an “extrinsic” theory that necessitates embedding in external coordinate systems for its description. From this perspective, contemporary quantum theory can only be seen as a version of “Newtonian mechanics” built upon the Cartesian coordinate system. This is also one of the key reasons for the conflict between the framework of quantum theory and that of the general principle of relativity. Such an intrinsic, background-independent quantum framework is particularly urgent for understanding the quantum properties of gravity based on the intrinsic spacetime geometry. Quantum theory and General Relativity must be established on a unified conceptual foundation.
Since the traditional quantum theory necessitates dividing the world into two distinct parts—a quantum system under-study and an external absolute classical apparatus or classical “observer”—it requires the introduction of several highly controversial assumptions to provide a coherent explanation of the contemporary quantum mechanics. For instance, according to the standard Copenhagen interpretation, a quantum wavefunction with a large spatial extent and extension must undergo an instantaneous “superluminal” collapse across two distant spatial coordinates (as exemplified by the so-called EPR paradox). Einstein used such a paradox to express his perceived intrinsic incompatibility between quantum theory and the relativity. Current technological advancements have enabled the conduct of long-distance quantum mechanics experiments spanning several kilometers or even hundreds of kilometers (such as experiments involving the distribution and measurement of entangled photons at long distances apart). Although the correlation measurement results of the entangled photon pairs agree to the predictions of the contemporary quantum mechanics, the issue does not lie with the mathematical formalism of quantum mechanics itself, but rather with its physical interpretation. Specifically, it concerns how an observer at one spatial location can locally prepare experimental apparatus to measure a local quantum state and how to appropriately interpret the “relative relationship” (rather than the two individual quantum states themselves) between this measured quantum state and a quantum state at another distant spatial location. Currently, there is no superior interpretation of the mathematical formalism of the contemporary quantum mechanics that provides a self-consistent and unified framework for understanding these inherent contradictions. The Copenhagen interpretation remains the closest to a “laboratory-common-sense” and a most convenient explanation for the mathematical formalism of contemporary quantum theory, leading to its widespread acceptance among physicists working at laboratory scales. However, we recognize that when certain interpretations (through thought experiments) are pushed to extreme scenarios, they may lead to inherent problems.
Due to the aforementioned shortcomings of the contemporary quantum mechanics, this paper attempts to establish a quantum mechanical framework that is (1) intrinsic, background-independent, and geometrized—that is, it relies solely on the description of internal physical degrees of freedom without resorting to any external absolute parameters; (2) treating both the quantum system under-study and the quantum measuring apparatus on an equal footing; and (3) formulates a theory that describes the “interrelationships” between quantum states, particularly the relationships between the quantum state of the system under-study and that of the measuring quantum apparatus, rather than focusing on individual quantum states relative to an external absolute classical apparatus. In other words, it employs relative or conditional probabilities to replace the standard absolute probabilities. In this paper, we will observe that entangled states provide an appropriate conceptual foundation for “internalizing” or “relationalizing” the interpretation of quantum mechanics, rather than serving as absolute quantum states in the framework of the contemporary quantum mechanics. This approach thereby avoids many of the perplexing issues arising from the so-called “external” interpretations of the standard Copenhagen interpretation.
Because the vast body of literature on the subject on the relational formulations and internal evolution of quantum mechanics, it is hardly for us to fully grasp all the existing works to review here, we only provide a potentially incomplete overview of some key landmark studies that we consider crucial and relevant to the topic of this paper, and how our works differ from these aforementioned studies are leaving in the section-VII. The first internal clock time evolution formalism of quantum mechanics can be attributed to the seminal work of Page and Wootters (PhysRevD.27.2885, ; 1986Quantum, ) in 1980s. And there are also a large subsequent literature on the Page-Wootters mechanism, relational time and quantum clock, see e.g. (Foti:2020erm, ; Woods2019pagewootters, ; qfns-48vq, ) and references therein. The concept of quantum reference frame is also a long-standing research in literature, including early foundational works (PhysRevD.30.368, ) and extensive developments on superselection rules (RevModPhys.79.555, ), relational observables (Rovelli:1990pi, ), frame transformations (Giacomini:2017zju, ), perspective dependence (2020Quant…4..225V, ; Hohn_2020, ) and symmetry-based formulations (Loveridge:2017pcv, ) and large number of relevant references therein. And the quantum reference frame relies on a relational interpretation of quantum physics, which started from Rovelli’s seminal paper (Rovelli:1995fv, ). On the other hand, through analogies between the Hilbert space and the Riemannian spaces endowed with additional geometric structures (such as complex and symplectic structures), numerous attempts have been made to geometrize quantum mechanics and incorporate it into the realm of pure geometry (see (Kibble:1978tm, ; Provost:1980nc, ; Ashtekar:1997ud, ; Brody:1999cw, ; Grabowski:2005my, ; Bengtsson:2006rfv, ) and the references therein). Among these, (Pandya:2006mrv, ; Matassa:2020dnc, ) place particular emphasis on the application of the Fubini-Study metric, a special metric in the Complex Projective Space, in describing quantum states, as well as the geometry underlying non-dynamical phases of quantum states (Simon:1983mh, ; Page:1987ca, ; Anandan:1990fq, ; Artacho:2017zey, ). Meanwhile, (Mosseri:2001gbm, ; Levay:2003pb, ; Grabowski:2005my, ) focus on understanding quantum entangled states from a geometric perspective. There have been many endeavors to establish an intrinsic, background-independent, and geometrized quantum mechanics. (Minic:2003en, ; Minic:2003nx, ) propose a quantum evolution equation akin to Einstein’s equation and provide a geometrized interpretation of probabilities. The problem of time in covariant quantum gravity is given by e.g. (Rovelli:1995fv, ; Gambini:2006ph, ), and the conditional probability interpretations of the Hamiltonian constraint is given by e.g. (Dolby:2004ak, ).
This paper takes a quantum clock as an example of the simplest reference system to explore how the motion of object’s position and the readings of the quantum clock are incorporated into a fundamental description of quantum mechanics. Thus, the quantum mechanics discussed in this paper is of a specific type. When compared with the functional integral approach (Luo:2021zpi, ; 2024arXiv240809630L, ), this method based on entangled states or relative states is equivalent, yet each has its own advantages. The relative state approach discussed in the paper faces difficulties in being generalized to scenarios involving quantum spacetime reference frames (Luo2014The, ; Luo2015Dark, ; Luo:2015pca, ; Luo:2019iby, ; Luo:2021zpi, ; Luo:2022goc, ; Luo:2022statistics, ; Luo:2022ywl, ; Luo:2023eqf, ; 2023AnPhy.45869452L, ; 2024arXiv240809630L, ; Luo:2025cer, ) and quantum field theory, whereas the functional integral approach is relatively straightforward in this regard. The functional method also more readily reveals the relationship between inertial forces and gravity that arises from general coordinate systems. However, this canonical quantum mechanical description based on relative states is more suitable for conceptually demonstrating its connections to the standard Schrödinger equation in quantum mechanics and to the standard absolute probability interpretation of the contemporary quantum mechanics than the functional integral approach, which is the main motivation of the paper.
In Section II of this paper, we first examine, from a historical perspective, how clocks, as a standard reference system, have been introduced and evolved in physics. In Section III, we discuss how to interpret the relative relationship between the quantum state of object’s coordinates and the quantum state of a clock from the perspective of the entangled states. We also explore the general relative interpretation of the entangled states and its underlying geometry: the non-trivial fiber bundles. In Section IV, from a purely geometric standpoint, we discuss how the quantum state of object’s coordinates evolves relative to the quantum state of the clock. In Section V, the relationship between the derived evolution equation and the Schrödinger equation is discussed. In Section VI, several non-inertial effects that arise in the reference frame of this quantum clock is discussed. We give some comparision between our work and the past works in Section VII. Finally, we conclude the paper with a summary. In the appendix, we review the functional integral approach, which also deals with quantum clocks, for the purpose of comparing it with the relative state approach.
II Clock Time as a Reference System
Before going into details how a clock time measuring the so-called “time”, let us first take a broad look at the concept of spacetime as a reference system based on the quantum equivalence principle (2024arXiv240809630L, ). In fact, we have never “truly” measured spacetime itself, all our measurements of the so-called spacetime are indirect, relying on some material reference system (such as light or other material fields serving as references). A quantum equivalence principle (2024arXiv240809630L, ) provides the physical foundation for establishing a quantum reference frame theory through the measurement of the spacetime using material reference systems. Based on this principle, we know how to use material measuring rods and clocks to measure those universal properties (independent of the mass and other properties of the material reference system) that can be reinterpreted as the geometric properties of the spacetime, rather than merely properties of the measuring rods and clocks themselves. These universal properties of the spacetime reflected in material reference systems can manifest not only at the classical level as first-order moment averages independent of the material reference system’s mass (such as the averaged coordinates) but also at the quantum level as second-order moment quantum fluctuations of the material reference system (such as the coordinate broadening). Attributing these universal properties of material reference systems to spacetime may stem from human’s priori instinct or a desire for conceptual simplicity, akin to how, in early human history, barter trading (“trading goods for goods”) was inconvenient when comparing the values of different items. Subsequently, the development of a universal currency, which abstracts and measures the general value in all transactions and serves as a general equivalent, greatly simplified trading. Similarly, the spacetime reference frame plays the role of a general equivalent for comparing different motions. Clock time, as the simplest example of a spacetime reference frame, is the focus of this paper. After clarifying the example of the clock time, the concept of spacetime reference frames can naturally be extended in a similar manner.
The establishment of Newton’s laws of motion is based on the ability to physically measure time using mechanical clocks. Galileo discovered the isochronism of pendulums, which could serve as the operational principle for mechanical clocks. Although there was no “standard clock” at that time to explain to Galileo why pendulums were “isochronous”, he could only rely on the only “standard” available to him: his pulse. By Newton’s era, the concept of time had been mathematized. Analogous to the parameterization of complex curves, Newton postulated the existence of a mechanical device, akin to Galileo’s pendulum, that automatically “generates” a parameter—(Newtonian) time—for the moving objects under study. When seemingly complex motions in nature are parameterized by this time variable, they are greatly simplified. To establish his laws of motion, Newton’s first step was to define a clock through his first law. Under this defined clock, the equation of Newton’s second law takes on the simplest form. Newton’s laws of motion aim to determine the forms of the spatial coordinates $X$, $Y$, and $Z$ of an object moving in three-dimensional space when parameterized by the Newtonian time $t$, denoted as $X(t)$, $Y(t)$, and $Z(t)$, which are referred to as the equations of motion for the object.
However, after people gradually gained an understanding of electromagnetism through experiments in the 19th century, the framework of the parameter—generated by Newton using a pendulum—came under challenge. This is because Maxwell’s equations in electromagnetism do not satisfy Galileo’s principle of relativity. When Maxwell’s equations are transformed into a uniformly moving coordinate system, their mathematical form changes. In other words, it seemed possible to determine, through electromagnetic experiments, whether an observer was undergoing absolute motion, which contradicted the relativity principle that motion is only relative. So, was it Maxwell’s equations that were incorrect in moving coordinate systems, or was it Galileo’s principle of relativity that was flawed? It was not until Einstein that the crux of the problem became clear: to ensure that Maxwell’s equations and the principle of relativity both hold true in moving coordinate systems, Newton’s notion of the global parameter $t$ representing time had to be abandoned. In Einstein’s view, time is the value $T$ that an observer reads from his/her clock, rather than Newton’s parameter $t$. Moreover, the clock reading $T$ that an observer obtains varies across coordinate systems moving at different velocities, provided that the speed of light is a constant in all moving frames. Thus, to describe the motion of an object, in addition to the three spatial coordinates $X(\tau)$, $Y(\tau)$, and $Z(\tau)$, there must be also Einstein’s clock reading $T(\tau)$, assuming the existence of a global parameter $\tau$. This hypothetical global parameter $\tau$ merely serves as a common parameter for $X(\tau)$, $Y(\tau)$, $Z(\tau)$, and $T(\tau)$, and does not necessarily have to be related to time any more, although we often refer to it as the “proper time” for historical reasons. Einstein discarded Newton’s unobservable global parameter $t$ as time and instead used the physical clock readings $T$ in coordinate systems moving at different velocities as an intermediate equivalent for comparing different motions within the same inertial frame: (Einsteinian) time. Consequently, the equations of motion for objects in inertial frames moving at different velocities transformed from a set of functions $X(t)$, $Y(t)$, and $Z(t)$ into a set of functionals $X[T(\tau)]$, $Y[T(\tau)]$, and $Z[T(\tau)]$.
Although Einstein redefined time in a stricter sense within physical systems, Newton’s global parameter $t$ still persists in the contemporary quantum theory presented in current textbooks. In quantum mechanics, Heisenberg was the first to abandon the concept of unobservable electron orbits within atoms, instead focusing solely on observable quantities, such as atomic spectral lines. In Bohr’s early interpretations, since each spectral line arises from transitions between two electron energy states, so in Heisenberg’s framework, the electron’s spatial coordinates cease to be ordinary classical numbers (c-numbers) and instead become matrices (q-numbers) representing transitions between different energy states. In other words, the electron’s spatial coordinates transform from a single number into a matrix. Matrix mechanics marked the first step towards the textbook quantum mechanics we recognize today. Although the spatial coordinates of microscopic particles are re-interpreted in quantum mechanics, time remains Newton’s global parameter. To ensure that quantum theory satisfies Lorentz invariance and holds true in Lorentz inertial frames, quantum field theory merely replaces Newton’s single global parameter $t$ with four global parameters $x,y,z,t$ interpreting them as Minkowski spacetime coordinates. Quantum mechanics presupposes a classical external observer wielding Newton’s parameter $t$ generator, while quantum field theory’s classical external observer generates four global parameters. The existence of such external observers leads to fundamental difficulties in the contemporary quantum theory. For instance, it cannot be applied to the entire universe because, by definition, the universe has no exterior. Moreover, quantum theory’s division of the world into an observed quantum system and an external classical observer/instrument necessitates additional assumptions in contemporary quantum theory to handle the measurement process between the external classical observer and the observed quantum system, such as the collapse of wave function in the Copenhagen interpretation.
Einstein’s theory of relativity informs us that time is no longer Newton’s global external parameter $t$, but rather the clock reading $T(\tau)$ in each inertial frame. A clock is a physical instrument that “generates” a standardized motion $T(\tau)$ as an idealized reference, such as the uniform motion of an object described by Newton’s first law or the uniform circular motion of a clock’s pointer. However, we know from quantum theory that all motions in the world are non-idealized because they are always subject to quantum fluctuations. In other words, there may not exist an absolutely idealized, fluctuation-free quantum clock in the world. Nevertheless, the fact that an object serving as a clock is always subject to quantum fluctuations may not be of great significance, since the clock merely acts as an intermediate equivalent or reference point to facilitate comparisons between other motions, thereby simplifying complex motions. It is analogous to how, when exchanging goods directly, the values of the goods can be complex and difficult to discern. However, when a common intermediate equivalent, “money”, is used for exchanges, the values of the goods become immediately apparent. The fluctuations in the value of the “money” itself are not important, what matters is the relationship between different goods and different motions.
To mathematically define a quantum clock as a reference system for comparing different motions, the first question that arises is how to quantize Einstein’s clock $T(\tau)$ in such a way that it not only adheres to quantum mechanics, experiencing quantum fluctuations, but also satisfies the principle of relativity. Additionally, the second question we must pose is: what is the precise meaning, at the quantum level, of the classical-level functional equations of motion $X[T(\tau)]$, $Y[T(\tau)]$, and $Z[T(\tau)]$?
III Entangled State as a relative state
For simplicity, let’s consider a one-dimensional moving object with a spatial coordinate $X(\tau)$, whose motion with respect to the parameter $\tau$ is governed by a Hamiltonian $H_{X}$. Additionally, we now introduce a clock, with its pointer coordinate $T(\tau)$ evolving with respect to $\tau$ as described by a Hamiltonian $H_{T}$. Here, the global parameter $\tau$ shared by both $X$ and $T$ can be interpreted as a global Newtonian parameter. Furthermore, we assume that there is no interaction between the object $X(\tau)$ and the clock $T(\tau)$. They evolve independently with respect to the parameter $\tau$. Thus, the Hamiltonian of the entire system is given by $H=H_{X}+H_{T}$. The quantum state space of the entire system is the direct product space
$\mathscr{H}_{X}\otimes\mathscr{H}_{T}$ (1)
where $\mathscr{H}_{X}$ is the quantum state space spanned by the object’s coordinate $X(\tau)$ and $\mathscr{H}_{T}$ is the quantum state space spanned by the clock’s pointer coordinate $T(\tau)$. However, the quantum state of the entire system, as an eigenstate of the total Hamiltonian, is not necessarily a simple direct product state of the object’s quantum state $|X(\tau)\rangle$ and the clock’s quantum state $|T(\tau)\rangle$. Generally, it is an entangled state, with a dimensionality $N$ equal to that of the smaller Hilbert subspace
$|X,T\rangle=\sum_{\tau}^{N}C_{\tau}|X(\tau)\rangle\otimes|T(\tau)\rangle$ (2)
All measurement processes involve a comparison between the state of the system under-study and the state of the measuring instrument. Prior to conducting a measurement on the system, an initialization step known as instrument calibration (or scale mark) is necessary. This step involves adjusting the states of the object’s coordinate $X$ and the clock $T$ to establish a one-to-one correspondence between the instrument (the clock $T$) and the system under-study (the object position $X$), thereby enabling the measurement of the system’s state by reading the instrument’s state. The “one-to-one correspondence” is equivalent to being able to extract the maximum information from the under-study system. The superposition coefficients $C_{\tau}$ of the entangled state (2) are determined by the state preparation during calibration. For instance, the system can be adjusted so that at $\tau=0$, the clock pointer is initialized in the state $T(0)=0$, while the object’s coordinate is initialized at the position $X(0)=X_{0}$. This process is equivalent to subjecting them to an instantaneous interaction, after which they cease to interact and instead evolve independently according to their respective Hamiltonians with respect to $\tau$. During subsequent evolution, whenever the clock is in a certain evolved state $|T(\tau)\rangle$, the object is also necessarily in a corresponding evolved state $|X(\tau)\rangle$. Thus, the state of the entire system is at $|X(\tau)\rangle\otimes|T(\tau)\rangle$, related to an amplitude of $C_{\tau}$ in front and a probability of $|C_{\tau}|^{2}$. The amplitude or probability that the object is in a corresponding state when the clock is in a certain state is prepared during instrument calibration. Moreover, quantum principles permit the existence of superposition of these one-to-one states, where all possible one-to-one states of the clock and object position are superposed with the corresponding amplitudes $C_{\tau}$, resulting in the entangled state (2). From a mathematical perspective, this entangled state establishes a one-to-one (probabilistic) mapping $|X,T\rangle:|T(\tau)\rangle\rightarrow|X(\tau)\rangle$ between the object’s position states $|X(\tau)\rangle$ under possible clock states $|T(\tau)\rangle$. This mathematical mapping is essentially the counterpart of the functional mapping $X[T(\tau)]:T(\tau)\rightarrow X(\tau)$, so in this sense, the entangled state can be regarded as a quantum version of the functional equation of motion $X[T(\tau)]$.
Unlike deterministic events described by functional equations of motion $X[T(\tau)]$, this entangled state depicts a probabilistic event. When the clock is in the state $|T(\tau)\rangle$, the object’s position is in the state $|X(\tau)\rangle$. The normalized joint probability of these two events occurring simultaneously is $|C_{\tau}|^{2}$ (w.r.t. an external observer). Unlike a simple direct product state $|T\rangle\otimes|X\rangle$, the entangled state $|X,T\rangle$ is inseparable, meaning that the joint probability $|C_{\tau}|^{2}$ does not equal the simple product of the normalized probability $|A_{\tau}|^{2}$ of measuring the clock alone in the state $|T(\tau)\rangle$ and the normalized probability $|B_{\tau}|^{2}$ of measuring the object’s position alone in the state $|X(\tau)\rangle$, i.e. $|C_{\tau}|^{2}\neq|A_{\tau}|^{2}|B_{\tau}|^{2}$. Here, $|T\rangle=\sum_{\tau}A_{\tau}|T(\tau)\rangle$ and $|X\rangle=\sum_{\tau}B_{\tau}|X(\tau)\rangle$ represent certain quantum states that each subsystem alone would be in, expanded in terms of their respective basis vectors in the basis of the entangled state. Therefore, the probability of the object’s position being in the state $|X(\tau)\rangle$ given that the clock is definitely in the state $|T\rangle$ is a conditional probability (or relative probability), which can be derived using the conditional probability formula
$\textrm{P}\left(X(\tau)|T\right)=\frac{\textrm{P}\left(X(\tau)\cup T(\tau)\right)}{\textrm{P}\left(T(\tau)\right)}=\frac{\textrm{P}\left(X(\tau)\cup T(\tau)\right)}{\textrm{P}\left(T(\tau)|T\right)}=\frac{|C_{\tau}|^{2}}{|A_{\tau}|^{2}}.$ (3)
where $\textrm{P}\left(T(\tau)\right)=\textrm{P}\left(T(\tau)|T\right)=|A_{\tau}|^{2}$ represents the probability (w.r.t. an external observer) that the reference state $|T(\tau)\rangle$ is found in the subsystem state $|T\rangle$ of the entangled state, while $\textrm{P}\left(X(\tau)\cup T(\tau)\right)=|C_{\tau}|^{2}$ denotes the joint probability (w.r.t. an external observer) that the entangled state is simultaneously in the states $|X(\tau)\rangle$ and $|T(\tau)\rangle$. However, their ratio, the relative probability is then independent to the external observer.
We refer to the quantum state $|T\rangle$ of the quantum clock or measuring instrument as the conditional state (or reference state). It is input by observer and depends on the specific conditions under which the observer wishes to obtain the relative probabilities of the entangled state $|X,T\rangle$ as output. This constitutes the “relative probability” interpretation of the entangled states. It is a distinct interpretation from the “absolute probability” interpretation in the standard quantum mechanics, as it additionally takes into account the conditions of the state occupied by the instrument’s subspace. Within the framework, to extract physically meaningful measurement results from the entangled state that describes the relative relationship between the under-study quantum system and the quantum instrument, one must rely on the relative probability interpretation.
We observe that the entangled state $|X,T\rangle$ cannot be expressed as a direct product state, i.e. $|X,T\rangle\neq|X\rangle\otimes|T\rangle$. The relative probability of the object’s position being in the state $|X(\tau)\rangle$ given that the clock is in the state $|T\rangle$ does not equal the probability obtained from a separate measurement of the object’s position, i.e. $\textrm{P}\left(X(\tau)|T\right)=\frac{|C_{\tau}|^{2}}{|A_{\tau}|^{2}}\neq|B_{\tau}|^{2}$. Therefore, the inseparable entangled state $|X,T\rangle$ can only be interpreted probabilistically based on the relative relationship between the two states. The state of the object’s position $|X(\tau)\rangle$ only holds physical meaning relative to the state of the clock $|T(\tau)\rangle$. Discussing the state of the object’s position alone or the state of the clock alone lacks absolute meaning, as their exhibited quantum probabilistic behaviors are entirely different.
The inseparability of the entangled state (2), or its inability to be expressed as a global direct product state, mathematically means that the entangled state $|X,T\rangle$ cannot be represented by a single global clock state that encompasses all possible position states. Instead, it can only locally map out a position state $|X(\tau)\rangle$ on a local clock state $|T(\tau)\rangle$, and then superpose (or “glue together”) the states of all local bases to obtain the overall quantum entangled state. In other words, the entangled state $|X,T\rangle=\sum_{\tau}C_{\tau}|X(\tau)\rangle\otimes|T(\tau)\rangle$ forms a non-trivial fiber bundle with a fiber $|X(\tau)\rangle$ growing over the local base space $|T(\tau)\rangle$, and a projection $T^{-1}$
$T^{-1}:|X,T\rangle\rightarrow|X(\tau)\rangle$ (4)
from the fiber bundle $|X,T\rangle$ to a section $|X(\tau)\rangle$ of the local fiber, gives rise to the relative amplitude (see below). The fiber bundle $|X,T\rangle$ is non-trivial means that it is not equivalent to a globally direct product state $|X\rangle\otimes|T\rangle$, i.e. the entangled state $|X,T\rangle$ corresponds to a non-trivial fiber bundle. Compared with that, in the standard quantum mechanics, the wave function $|X(t)\rangle$ is trivially a global section of the fiber over a global Newtonian time $t$ as the base space, meaning that it is a trivial fiber bundle.
Only when the probability of the object’s position state $\textrm{P}\left(X(\tau)|T\right)$, equals the probability obtained from separate measurements of the object’s position and the clock, i.e., when the state is separable i.e. $C_{\tau}|^{2}=|A_{\tau}|^{2}|B_{\tau}|^{2}$, can we use a single global time to cover all possible position states. In this case, the entangled state $|X,T\rangle$ reverts to the separable direct product state $|X\rangle\otimes|T\rangle$ or a trivial fiber bundle, allowing the entire background state of the clock to be separated out and reduced to the standard textbook quantum state $|X\rangle$, which describes the probability $|B_{\tau}|^{2}$ of measuring the object’s position $X$ at a global time $\tau$ ($|B_{\tau}|^{2}$ has physical meaning w.r.t. $\tau$). Therefore, we require a relativistic quantum mechanical interpretation based on the “relative relationship between the two states $|X(\tau)\rangle$ and $|T(\tau)\rangle$” described by the local entangled states, to generalize the Copenhagen interpretation of quantum mechanics for “absolute quantum states $|X(\tau)\rangle$” in textbooks. This is analogous to generalizing global Cartesian geometry to curved Riemannian geometry based on tangent bundles over local base spaces. Establishing a relativistic quantum theory and an intrinsic interpretation that measures one entangled subspace (the quantum system under-study) relative to another subspace (the quantum measuring instrument) within a larger quantum state space is particularly urgent in explaining a quantum theory of gravity.
An important consequence of the distinction between the non-trivial fiber bundle described by the entangled states and the global trivial fiber bundle depicted in conventional quantum mechanics lies in the fact that, due to the absence of a global base space in the non-trivial fiber bundle, the local basis vectors on the base space $T$ are not necessarily orthogonal (i.e. nonorthogonal). Instead, they define a non-trivial metric on the base space $T$ given by the inner product $\langle T(\tau)|T(\tau^{\prime})\rangle=s^{\tau\tau^{\prime}}$, which is not necessarily gauge equivalent to the Kronecker’s delta $\delta_{\tau\tau^{\prime}}$ of the flat case
$\langle T(\tau)|T(\tau^{\prime})\rangle=\langle T(\tau)|T\rangle\langle T|T(\tau^{\prime})\rangle=A^{\tau*}A^{\tau^{\prime}}=\frac{A^{\tau*}}{A_{\tau^{\prime}}}=A^{\tau}A^{\tau^{\prime}*}=\frac{A^{\tau}}{A_{\tau^{\prime}}^{*}}=s^{\tau\tau^{\prime}}=\left(s\right)_{\tau\tau^{\prime}}^{-1}\neq\delta_{\tau\tau^{\prime}}$ (5)
In other words, the base space of a non-trivial fiber bundle can be seem “curved”, unlike the “flat” basis vectors of the global trivial fiber bundle in conventional quantum mechanics. The orthonormality or invariant inner product must now be expressed in terms of the basis vectors $|T\rangle$ and their dual basis vectors $\langle\bar{T}|$
$\langle\bar{T}(\tau)|T(\tau^{\prime})\rangle=A_{\tau}^{*}A^{\tau^{\prime}}=\langle T(\tau)|\bar{T}(\tau^{\prime})\rangle=A_{\tau}A^{\tau^{\prime}*}=\delta_{\tau}^{\tau^{\prime}}$ (6)
As long as the expansion coefficients $A_{\tau}$ are normalized, it follows that $\sum_{\tau}A_{\tau}^{*}A_{\tau}=\sum_{\tau}s_{\tau\tau}=1$.
The second method for calculating the relative amplitude is by using the projection of the fiber bundle (4): the relative amplitude is obtained by performing a “conditional projection” of the vector $|X,T\rangle$ in the non-trivial fiber bundle onto the section $|X(\tau)\rangle$ of the fiber, where the reference state $|T\rangle$ as a projector can be expanded by using the basis vectors $|T(\tau)\rangle$ in the entangled state
$\langle T|$ $=\sum_{\tau}A_{\tau}^{*}\langle T(\tau)|=\sum_{\tau,\tau^{\prime}}A_{\tau}^{*}s^{\tau\tau^{\prime}}\langle\bar{T}(\tau^{\prime})|=\sum_{\tau,\tau^{\prime}}A_{\tau}^{*}\left(A^{\tau^{\prime}}A^{\tau*}\right)\langle\bar{T}(\tau^{\prime})|$
$=\sum_{\tau,\tau^{\prime}}\delta_{\tau}^{\tau^{\prime}}A^{\tau*}\langle\bar{T}(\tau^{\prime})|=\sum_{\tau}A^{\tau^{\prime}*}\langle\bar{T}(\tau^{\prime})|=\sum_{\tau^{\prime}}\frac{1}{A_{\tau^{\prime}}}\langle\bar{T}(\tau^{\prime})|$ (7)
By using the conditional projection by $\langle T|$, we have a projected entangled state
$\langle T|X,T\rangle=\sum_{\tau}C_{\tau}|X(\tau)\rangle\langle T|T(\tau)\rangle=\sum_{\tau,\tau^{\prime}}C_{\tau}A_{\tau}^{*}\langle T(\tau^{\prime})|T(\tau)\rangle|X(\tau)\rangle=\sum_{\tau,\tau^{\prime}}C_{\tau}A_{\tau}^{*}s^{\tau\tau^{\prime}}|X(\tau)\rangle=\sum_{\tau}\frac{C_{\tau}}{A_{\tau}}|X(\tau)\rangle$ (8)
in which the coefficient $\frac{C_{\tau}}{A_{\tau}}$ describes the relative amplitude (or conditional amplitude) of measuring the object’s position in the state $|X(\tau)\rangle$ given that the clock is in the state $|T\rangle$, reflecting the relative amplitude between the subsystems, then yielding the relative probability $\frac{|C_{\tau}|^{2}}{|A_{\tau}|^{2}}$.
Now, the magnitude (or the length) of the component $C_{\tau}$ of the state vector in the space $\mathscr{H}_{X}\otimes\mathscr{H}_{T}$ no longer holds absolute meaning (relative to the absolute normalization factor $\sum_{\tau}|C_{\tau}|^{2}$ and external parameter), instead, it only has relative meaning with respect to the magnitude $|A_{\tau}|^{2}$ of the component $A_{\tau}$ in the subspace $\mathscr{H}_{T}$. More precisely, what we refer to in this paper as describing the “relative relationship” between two mutually entangled systems essentially involves characterizing the “relative component” of the entangled state vector in one subspace relative to the other subspace. This “relative component” is precisely the relative amplitude or conditional amplitude.
The third method for calculating the relative amplitude or relative probability involves noting that this relative probability is also equivalent to taking the partial trace of the density matrix $\rho_{X,T}\equiv|X,T\rangle\langle X,T|$ of the entangled state with respect to the reference state $|T\rangle$
$\textrm{Tr}_{T}\left(\rho_{X,T}\right)=\langle T|X,T\rangle\langle X,T|T\rangle=\sum_{\tau,\tau^{\prime}}\frac{C_{\tau}}{A_{\tau}}\frac{C_{\tau^{\prime}}^{*}}{A_{\tau^{\prime}}^{*}}|X(\tau)\rangle\langle X(\tau^{\prime})|=\sum_{\tau}\left|\frac{C_{\tau}}{A_{\tau}}\right|^{2}|X(\tau)\rangle\langle X(\tau)|=\sum_{\tau,\tau^{\prime}}C_{\tau}C_{\tau^{\prime}}^{*}s^{\tau\tau^{\prime}}|X(\tau)\rangle\langle X(\tau^{\prime})|$ (9)
Here, taking the partial trace of the entangled state’s density matrix is not simply summing the diagonal elements within a subspace, but rather contracting by the metric $s^{\tau\tau^{\prime}}$of the subspace $T$, i.e. $\textrm{Tr}_{T}\left(\mathbf{O}\right)=\mathbf{O}_{\mu\nu}s^{\mu\nu}\neq\mathbf{O}_{\mu\nu}\delta_{\mu\nu}$. This is analogous to how, in a curved Riemannian space, taking the trace of a matrix is no longer a straightforward summation of all diagonal elements, instead, it requires contraction using the curved metric. This represents a significant difference between quantum mechanics based on entangled states and conventional (global) quantum mechanics.
Because the amplitude of the entangled state, $C_{\tau}$, is determined by the calibration and/or preparation of the instrument, rather than by the fundamental equations of theory, and remains fixed during the experimental process, so the entangled state merely serves as a “cross-reference table” between the instrument readings and the state of the system under-study prepared during the calibration of the system under-study and the quantum instrument. The relative amplitude obtained through metric projection onto the conditional state, as well as its evolution (relative to the instrument subsystem), are entirely determined by the metric of one subsystem $X$ and its evolution relative to another subsystem $T$ of the entangled state. The relative probability given by the relative amplitude is physically measurable. The evolution of the metric of the subspace $X$ relative to $T$ is coming from the Ricci-flat Kähler-Einstein equation (see the next chapter). This equation links the intrinsic curvature of the metric of the non-flat subspace $\mathscr{H}_{X}$ embedded in the flat quantum state space $\mathscr{H}_{X}\otimes\mathscr{H}_{T}$ with the extrinsic curvature relative to the evolution of the another subspace $\mathscr{H}_{T}$. It serves as the equation of motion for the metric of one subspace $\mathscr{H}_{X}$ relative to the state vectors of the other subspace $\mathscr{H}_{T}$. Similarly, we also have the evolution equation for the metric of the subspace $\mathscr{H}_{T}$ relative to $\mathscr{H}_{X}$, which describes its relative evolution with respect to the state vectors of the subspace $\mathscr{H}_{X}$. The system under-study and the measuring instrument now are completely on an equal footing. More discussion on the relationship between evolution equation and its degradation to the Schrödinger equation will be provided in the following paper. Thus, we observe that the evolution of the relative amplitude is not generated by the evolution of the entangled state (which remain fixed after the calibration), but rather by the relative evolution of the subspace metrics or basis vectors of the Hilbert subspaces relative to another subspace. This differs from standard quantum mechanics, where the initially prepared quantum state undergoes unitary evolution (unitary rotation) relative to external parameters. Here, the entangled state vector does not evolve relative to external parameters, instead, relative evolution occurs between Hilbert subspaces, causing the components of the entangled state $C_{\tau}$ projected onto the subspace $\frac{C_{\tau}}{A_{\tau}}$ (or $\frac{C_{\tau}}{B_{\tau}}$) to evolve relative to another subsystem. Adopting the form of the evolution equation for the intrinsic representation of quantum states allows us to treat the system under-study (object’s position $X$) and the measuring instrument (physical clock $T$) on an equal footing and symmetrically, in other words, we can also treat the physical clock as the system under-study and the object’s position as the measuring instrument.
The derivative of the basis vector $|T(\tau)\rangle$ of the base space with respect to $\tau$, namely $\langle T(\tau)|\frac{d}{d\tau}|T(\tau)\rangle$, yields a non-trivial connection. This non-trivial connection on the curved base space leads to the emergence of a non-integrable Berry phase $\Delta\varTheta$ when the state vector undergoes parallel transport across the base space
$\Delta\varTheta=\int_{0}^{\hat{\tau}}d\tau\langle T(\tau)|\frac{d}{d\tau}|T(\tau)\rangle=\oint_{c}d\mathbf{x}(\tau)\cdot\langle T[\mathbf{x}(\tau)]|\mathbf{\nabla}_{\mathbf{x}}|T[\mathbf{x}(\tau)]\rangle=\oint_{c}d\mathbf{x}(\tau)\cdot\mathbf{K}[\mathbf{x}(\tau)]$ (10)
The non-vanishing of the phase change along a closed loop means that there is no longer a global direct product of state vector, the state formed by these two subsystems must be an entangled state. (To facilitate the observation of the relationship with the standard Berry phase, we can define $\mathbf{x}(\tau)$ as a higher-dimensional parameter space corresponds to $\tau$ . The connection of the state vector as the parameters change is $\langle T[\mathbf{x}(\tau)]|\mathbf{\nabla}_{\mathbf{x}}|T[\mathbf{x}(\tau)]\rangle\equiv\mathbf{K}[\mathbf{x}(\tau)]$. Starting from $\mathbf{x}(0)$ and traversing a closed loop $c$ in the higher-dimensional parameter space, we return to $\mathbf{x}(\hat{\tau})=\mathbf{x}(0)$. Here, $\mathbf{B}=\nabla_{\mathbf{x}}\times\mathbf{K}[\mathbf{x}(\tau)]$ represents the non-zero Berry curvature in the $\mathscr{H}_{T}$ subspace). Therefore, we observe that, in general, the inherent nature of non-trivial fiber bundles in the subsystem of an entangled state ensures the existence of a non-trivial Berry phase in the subspace and the subspace must be curved, in other words, the first Chern class or Chern number is non-trivial in the base space of the non-trivial fiber bundles.
The entangled state between the under-study system and that of a quantum instrument, along with its interpretation in terms of relative probabilities, provides the conceptual foundation for a quantum theory formulated in a general spacetime coordinate system. Entangled states are well-defined within the framework of quantum theory, yet they exhibit many properties that appear strange from the perspective of classical physics. In the early years of the quantum mechanics, it was assumed that the scale of the wave function was limited to atomic scale, and thus there was no need to worry about the issue of simultaneity arising from the instantaneous collapse of the wave function in the Copenhagen interpretation. However, Einstein’s EPR thought experiment pointed out that the scale of the wave function could, in fact, be much larger—so large that the wave functions are completely macroscopically orthogonal in space. This leads to the perplexing problem of “superluminal” effects associated with the instantaneous collapse of the wave function. Here, we merely offer an alternative perspective on the entangled state, viewing them as a type of “relative state” that describes the “relationship” between two states. This approach makes the picture of entangled state less incomprehensible. Take, for example, the spin version of the EPR thought experiment. A pair of spin-polarized photons or electrons, after undergoing a brief interaction—such as when light passes through a nonlinear medium, emitting two oppositely moving polarized photons A and B, leading to an entangled state, ofter called, the EPR state
$|\textrm{EPR}\rangle=\frac{1}{\sqrt{2}}\left(|\mathbf{n}\rangle_{A}\otimes|-\mathbf{n}\rangle_{B}\pm|-\mathbf{n}\rangle_{A}\otimes|\mathbf{n}\rangle_{B}\right)$ (11)
where $\mathbf{n}$ and $\mathbf{n}$ represent a pair of opposite spin directions for the particles in space. We avoid the conventional EPR state notation $|\textrm{EPR}\rangle=\frac{1}{\sqrt{2}}\left(|\uparrow\rangle_{A}\otimes|\downarrow\rangle_{B}\pm|\downarrow\rangle_{A}\otimes|\uparrow\rangle_{B}\right)$ because it misleadingly suggests that the local states are definitively up $\uparrow$ or down $\downarrow$. In reality, when we locally measure the spin of particle A, we find that its spin $\mathbf{n}$ is completely random. Similarly, measuring the spin of particle B alone also yields a completely random result. Although the spin directions of A and B are individually random, they are always opposite to each other, a fact that can be verified through a joint measurement. This indicates that the individual spin quantum states of particles A or B do not have absolute meaning. What is physically meaningful is the relationship between the spins of the two particles. The EPR state describes only this relationship, not the local properties of the particles. Just as it is meaningless to discuss the absolute position of an object without a reference frame or coordinate system, what is physically meaningful is the object’s position relative to the origin of a coordinate system. Therefore, our theory aims to describe not the spin states of individual local particles but rather the relationship between the spin states of the two particles. From this perspective, there is no need to worry about whether the collapse of the entangled state involves mysterious “superluminal” effects, because what “collapses” is a certain relationship, not local quantum properties. (In this sense, the interpretation of quantum collapse probability here is somewhat closer to the Bayesian notion of “a measurement of the degree of belief” as the “relationship” is more akin to an observer’s constructed belief about the objective reality. However, a “relationship” is of course not entirely subjective, the relationship between the two is an objective existence.) From the standpoint of the standard quantum state space of the overall entangled state, the “interrelationship” between the spin states of the two particles does indeed appear to collapse instantaneously to an external observer, just as the standard Copenhagen interpretation suggests. This does not violate relativity because there is no absolute external observer who can observe the “collapse” of the relative state from the outside. When the entangled state collapses to certain its component, the relationship between the two particles becomes deterministic for the “external observer” but the specific local individual spin of A or B (for the external observer) remain undetermined. In other words, this interrelationship does not inform observers at A and B about the absolute spin directions of their respective particles. The spin quantum states of A and B still require specific local measurements by observers at each location. For example, after A measures its spin state, A communicates this locally measured result to observer B via a signal that travels at most at the speed of light. Given A’s measurement result, observer B can then calculate the relative probability and expectation value for the direction of B’s measured spin state. Otherwise, B’s local observer perceives the spin direction as completely random relative to B’s local measurement apparatus. In other words, we can only measure the relative relationship between A and B, while the absolute states of A and B locally are completely random. This is because if you project the spin state prepared locally by a certain instrument, say $|\mathbf{m}\rangle$ (e.g., via a Stern-Gerlach gradient magnetic field prepared in a specific direction), the local spin state $|\mathbf{n}\rangle$ can form an arbitrary angle with the prepared spin state $|\mathbf{m}\rangle$, resulting in an arbitrary inner product and relative phase. To measure the relative phase (or corresponding relative amplitude) between A and B, the two particles must be retransmitted at no more than the speed of light to the same point in space for an interference experiment. This interference experiment is equivalent to projecting the entangled state using the state of one of the subsystems as a reference state, yielding the relative amplitude or relative phase between them.
The relative or conditional probabilities derived from an entangled state reflect the fact that our information about the entangled system is incomplete in the sense that when one subsystem of the entangled state learns, via a classical channel, the measurement result of the other subsystem, it can utilize this additional information to update the prediction or the (posterior) probability about this local subsystem, as is claimed in the Bayesian theorem for an information incomplete system, such as the game of bridge where each player can only see a portion of the entire deck of cards.
In this EPR experiment, we can consider that the particle at location A serves as the measuring instrument or reference frame, while the other entangled particle B acts as the under-study system. The physically meaningful spin direction of the particle B is only relative to the spin direction of the measuring instrument, the particle A. Without knowledge of the state of the measuring instrument (the particle A), we cannot discuss the spin state of the particle B, and vice versa. An imperfect classical analogy would be the voltage difference between two widely separated electrodes. Suppose the voltage difference between them is fixed at 10V. This voltage difference can only have relative physical meaning. However, locally, the voltage measured at each electrode is entirely random, depending on the zero-point potential selected by the local observer’s measuring instrument. When you measure your local potential and find it to be, say, a completely random value aaV, and then communicate this value to another observer via a signal traveling at a finite speed, the potential of the other observer then “instantaneous” becomes (a+10)(a+10)V. This is because the voltage has only relative meaning, but in reality, there is no physical electric potential change experienced by the observer. Our quantum theory must also be entirely based on describing the relationships between different states, rather than on external absolute states. It should be founded on describing the relationship between the quantum states of the quantum system under-study and the quantum instrument, because, fundamentally, the measuring instrument is also part of the quantum world. It can exist in quantum superposition states and satisfy the quantum uncertainty principle, while the standard quantum theory assumes that the measuring instrument can only be in classical states and yield classical measurement results. Only when the measuring instrument can be approximately treated as a classical instrument, with its quantum effects ignored, do we revert to the textbook quantum theory that describes the quantum state of a single system being measured. Below, we will use a quantum clock as an simplest example to illustrate how to develop a dynamics of quantum theory upon describing the relationships between quantum states. We generalize this simple example to more general quantum spacetime reference frames to establish a quantum theory on more general spacetime coordinates and gravity (Luo2014The, ; Luo2015Dark, ; Luo:2015pca, ; Luo:2019iby, ; Luo:2021zpi, ; Luo:2022goc, ; Luo:2022statistics, ; Luo:2022ywl, ; Luo:2023eqf, ; 2023AnPhy.45869452L, ; 2024arXiv240809630L, ; Luo:2025cer, ).
IV Relative Evolution of States and Kähler-Einstein Equation
We will base our more detailed discussion on the relative evolution and “relative relationship” between the quantum state of the under-study system $|X_{i}\rangle\in\mathscr{H}_{X}$ and the quantum state of the measuring instrument or clock $|T_{a}\rangle\in\mathscr{H}_{T}$ on the fiber bundles or the geometry of the quantum state space $\mathscr{H}_{X}\otimes\mathscr{H}_{T}$ formed by the entangled state and its subspaces. This represents an intrinsic geometric perspective that treats the under-study subsystem and the quantum measuring instrument in a more symmetric manner.
To better highlight the similarities between the Hilbert space and the tensorial representation of Riemannian geometry, we will adopt a discrete basis vector formulation in this chapter, replacing the continuous basis vector version used in the previous chapter. Specifically, we will use $|X_{i}\rangle$ to replace $|X(\tau)\rangle$. Similarly, $|T_{a}\rangle$ represents the discrete version of $|T(\tau)\rangle$.
We start with the calibration process between the under-study system and the quantum instrument and prepares an entangled state
$|X,T\rangle=\sum_{i}^{N}C_{i}|X_{i}\rangle|T_{i}\rangle,\quad\left(C_{i}\in\mathbb{C}\right)$ (12)
This quantum state is a vector in the linear and flat quantum state space $\mathscr{H}_{X}\otimes\mathscr{H}_{T}$. To ensure a one-to-one correspondence between the quantum states of the two subspaces, the dimension of the total quantum state space is taken to be the dimension $N$ of the lower-dimensional subspace among the two. When this entangled state is collapsed in a local direct product basis $|X_{i}\rangle|T_{i}\rangle$, it gives a joint amplitude $C_{i}$ (it is generally a complex number) that satisfy the normalization condition $\sum_{i}^{N}|C_{i}|^{2}=1$.
It is worth noting that the prepared entangled state 12 is not a standard Schmidt decomposed entangled state. Here, $C_{i}$ is not Schmidt coefficient, and the local subspace basis vectors, which are selected from the calibration process, are not necessarily orthonormal. In contrast, the Schmidt decomposition requires selecting an appropriate local coordinate system $|X_{i}^{\prime}\rangle|T_{i}^{\prime}\rangle$ in the total quantum state space such that the expansion coefficients $C_{i}^{\prime}$ of the entangled state in this coordinate system are real-valued, and the local basis vectors are orthonormal.
Since a single subspace, either $\mathscr{H}_{X}$ or $\mathscr{H}_{T}$ may no longer be locally flat, we can always consider how the basis vectors $|X_{i}\rangle$ of the local $\mathscr{H}_{X}$ subspace vary with respect to the local quantum states in the local $\mathscr{H}_{T}$ subspace, whose state is expanded as
$|T\rangle=\sum_{a}^{N}t^{a}|T_{a}\rangle,\quad\left(t^{a}\in\mathbb{\mathbb{C}}\right)$ (13)
where $|T_{a}\rangle$ represents the (non-orthogonal) basis vectors in the local $\mathscr{H}_{T}$ subspace, and $t^{a}$ is the corresponding normalized amplitude. What we now aim to examine is the evolution of the basis vectors $|X_{i}\rangle$ in the local $\mathscr{H}_{X}$ subspace as they vary with the complex coordinates $t^{a}$ on the entangled local $\mathscr{H}_{T}$ subspace.
This entangled state constitutes a non-trivial fiber bundle structure formed by the bonding of local subsystems. Although the total quantum state space $\mathscr{H}_{X}\otimes\mathscr{H}_{T}$ is a flat linear space, its local subspace $\mathscr{H}_{X}$ is not flat. We consider the portion of the state space where $\mathscr{H}_{X}$ and $\mathscr{H}_{T}$ are entangled in a one-to-one correspondence, and thus its dimensionality is also $N$. The local metric can be defined through the inner product or overlap of the basis vectors $|X_{i}\rangle$
$h_{\bar{i}j}=\langle X_{i}|X_{j}\rangle\neq\delta_{ij},\quad h_{\bar{j}i}=\langle X_{j}|X_{i}\rangle=\langle X_{i}|X_{j}\rangle^{*}=h_{\bar{i}j}^{*}=h_{i\bar{j}}$ (14)
The bar over the lower index $i$ indicates that the bra vector corresponding to this lower index in the inner product has undergone a complex conjugation (since the physical metric is Hermitian, $h_{\bar{i}j}=h_{\bar{i}j}^{*}=h_{j\bar{i}}$, and the interchange of indices for this complex metric is symmetric). Similarly, the bar over the upper index $j$ denotes that the ket vector corresponding to the upper index in the inner product is a complex conjugate of the ket vector corresponding to the lower index
$h^{i\bar{j}}=\langle X^{i}|X^{j}\rangle,\quad h^{j\bar{i}}=\langle X^{j}|X^{i}\rangle=\langle X^{i}|X^{j}\rangle^{*}=h^{i\bar{j}*}=h^{\bar{i}j}$ (15)
Due to the Hermiticity of the metric, the metric components vanish when both indices are unbarred or both indices are barred. The basis vectors that are locally dual to the local basis vectors $|X_{i}\rangle$ are now denoted as $\langle X^{j}|$, and hence the bar in the bra vector of previous eq.(6) can be omitted in this convention (one upper index and one lower index). They satisfy the orthogonality and normalization conditions
$\langle X^{j}|X_{i}\rangle=\delta_{i}^{j},\quad\langle X_{j}|X^{i}\rangle=\delta_{\bar{j}}^{\bar{i}}\equiv\delta_{j}^{i}$ (16)
Now, $|X_{i}\rangle$ represents the basis vectors of the entangled state within the local subspace of the under-study system. Therefore, the quantum state $|X\rangle$ can also be expanded using this set of local basis vectors
$|X\rangle=\sum_{i}^{N}x^{i}|X_{i}\rangle,\quad\left(x^{i}\in\mathbb{\mathbb{C}}\right)$ (17)
in which $x^{i}$ represents the normalized complex coordinates.
Similarly, the metric of the subspace $\mathscr{H}_{T}$ can be given by the non-orthogonal basis vector
$s_{\bar{a}b}=\langle T_{a}|T_{b}\rangle\neq\delta_{ab},\quad s_{\bar{b}a}=\langle T_{b}|T_{a}\rangle=\langle T_{a}|T_{b}\rangle^{*}=s_{\bar{a}b}^{*}=s_{a\bar{b}}$ (18)
and
$s^{a\bar{b}}=\langle T^{a}|T^{b}\rangle,\quad s^{b\bar{a}}=\langle T^{b}|T^{a}\rangle=\langle T^{a}|T^{b}\rangle^{*}=s^{a\bar{b}*}=s^{\bar{a}b}$ (19)
$\langle T^{b}|$ is locally dual to $|T_{a}\rangle$ and $\langle T^{b}|$ is the discrete version of $\langle\bar{T}(\tau)|$ in eq.(6).
To derive the correct evolution equation, it is necessary to assume that the metric signature difference between $|X\rangle$ and $|T\rangle$ subspaces is opposite. In other words, we stipulate that the $|T\rangle$ subspace is timelike, meaning that it has a negative inner product
$\langle T^{b}|T_{a}\rangle=-\delta_{a}^{b},\quad\langle T_{b}|T^{a}\rangle=-\delta_{\bar{b}}^{\bar{a}}$ (20)
Meanwhile, $|X\rangle$ is spacelike and satisfies (16). The orthogonality condition (20) does not contradict the previously defined orthogonality in eq.(6) because we could just as well reverse the definitions entirely, assigning $|X\rangle$ a negative inner product and $|T\rangle$ a positive inner product instead. The roles of $|X\rangle$ as the under-study system and $|T\rangle$ as the measuring apparatus can also be symmetrically reversed, as long as the metric signature difference between $|X\rangle$ and $|T\rangle$ is opposite. This is because it is the relative signature difference between them that holds physical meaning (keeping the speed of light a constant).
Since locally each basis vector can be seen as a fiber of the complementary base subspace, the basis vectors of the two subspaces are locally orthogonal
$\langle T_{a}|X_{i}\rangle=0$ (21)
We can utilize the local metric $h_{\bar{i}j}$ of the subspace $\mathscr{H}_{X}$ to define the locally standard Levi-Civita connection and the intrinsic Ricci curvature on $\mathscr{H}_{X}$, the non-vanishing components are given by
$\Gamma_{ij}^{k}=h^{k\bar{l}}\frac{\partial h_{j\bar{l}}}{\partial x^{i}},\quad\Gamma_{\bar{i}\bar{j}}^{\bar{k}}=h^{\bar{k}l}\frac{\partial h_{\bar{j}l}}{\partial x^{i*}},\quad R_{\bar{i}j}=-\frac{\partial\Gamma_{\bar{i}\bar{k}}^{\bar{k}}}{\partial x^{j}}=-\frac{\partial}{\partial x^{j}}\left(h^{\bar{k}l}\frac{\partial h_{\bar{k}l}}{\partial x^{i*}}\right)$ (22)
in which the indices of $h_{\bar{i}j}$ are uppered or lowered by the metric of the same subspace, and we implicitly follow Einstein’s summation convention for tensors, summing over repeated upper and lower indices and omitting the summation symbols.
The evolution of the intrinsic metric $h_{\bar{i}j}$ of the subspace with respect to the “external” coordinates (time) outside the subspace, from a geometric perspective, arises from the concept of the extrinsic curvature of the subspace $\mathscr{H}_{X}$, describing how the hypersurface $\mathscr{H}_{X}$ is bending with respect to the external vector of the $\mathscr{H}_{T}$. By definition, the extrinsic curvature is the derivative of the intrinsic metric $h_{\bar{i}j}$ with respect to its normal “external” coordinate $t^{a}$
$K_{\bar{i}j}=\frac{1}{2}\frac{\partial h_{\bar{i}j}}{\partial t^{a}}|T^{a}\rangle$ (23)
which clearly describes the variation of the metric $h_{\bar{i}j}$ of the subspace $\mathscr{H}_{X}$ with respect to the state vector (13) of another entangled subspace $\mathscr{H}_{T}$. We can also define two types of scalar extrinsic curvature by contracting $K_{\bar{i}j}$ by $h^{\bar{i}j}$,
$K=h^{\bar{i}j}K_{\bar{i}j}=\frac{1}{2}h^{\bar{i}j}\frac{\partial h_{\bar{i}j}}{\partial t^{a}}|T^{a}\rangle,\quad K_{a}=\text{$\frac{1}{2}h^{\bar{i}j}\frac{\partial h_{\bar{i}j}}{\partial t^{a}}$}=\frac{1}{2}\frac{\partial}{\partial t^{a}}\left(\ln\det h_{\bar{i}j}\right)$ (24)
The derivative of the extrinsic curvature $K_{\bar{i}j}$ can be expressed as a second-order derivative of the metric $h_{\bar{i}j}$, from $\langle T_{b}|T^{a}\rangle=-\delta_{b}^{a}$, we have
$\frac{\partial K_{\bar{i}j}}{\partial t_{b}^{*}}\langle T_{b}|=\frac{1}{2}\frac{\partial^{2}h_{\bar{i}j}}{\partial t^{a}\partial t_{b}^{*}}\langle T_{b}|T^{a}\rangle=-\frac{1}{2}\frac{\partial^{2}h_{\bar{i}j}}{\partial t^{2}}$ (25)
Because the entangled state $|X,T\rangle$ is a standard quantum mechanical state, the entire Hilbert space $\mathscr{H}_{X}\otimes\mathscr{H}_{T}$ in which it resides is Ricci-flat, satisfying the Ricci-flat Kähler-Einstein equation
$R_{\bar{I}J}(q_{\bar{I}J})=0$ (26)
where $q_{\bar{I}J}=h_{\bar{I}J}+s_{\bar{I}J}$ is the subspace decomposition of the metric of the entire Hilbert space. When the Ricci-flat Kähler-Einstein equation is decomposed into the $\mathscr{H}_{X}$ and $\mathscr{H}_{T}$ subspace, besides encompassing $N$ constraint equations, our current focus lies on the $\frac{1}{2}N(N-1)$ relative evolution equations between the basis vectors of the subspaces. More precisely, we are interested in the equations that relate the intrinsic curvature $R_{\bar{i}j}$ of the $\mathscr{H}_{X}$ subspace to the extrinsic curvature $K_{\bar{i}j}$
$\frac{1}{2}\frac{\partial^{2}h_{\bar{i}j}}{\partial t^{2}}=R_{\bar{i}j}(h)+K(h)K_{\bar{i}j}(h)-2h^{k\bar{l}}K_{\bar{i}k}(h)K_{\bar{l}j}(h)$ (27)
which describes how the metric of $\mathscr{H}_{X}$ subspace evolves with respect to the external vector of the $\mathscr{H}_{T}$ subspace. And also analogous to the ADM decomposition of the Einstein equation, the equation can also be seen as complex version of the ADM decomposition of the Ricci-flat Kähler-Einstein equation. The geometric quantities (intrinsic curvature and extrinsic curvature) on both sides of the equation are covariant and no longer refer to any external absolute parameters. Therefore, this equation is generally covariant and provides an intrinsic description of the evolution of the quantum state basis vectors $|X_{i}\rangle$ (or the metric $h_{\bar{i}j}$) of the entangled subspace $\mathscr{H}_{X}$ relative to the state vector $|T\rangle$ (quantum clock) of the other entangled subspace $\mathscr{H}_{T}$.
Symmetrically, we can also have an equation for $s_{\bar{a}b}$
$\frac{1}{2}\frac{\partial^{2}s_{\bar{a}b}}{\partial x^{2}}=-\hat{R}_{\bar{a}b}(s)-\hat{K}(s)\hat{K}_{\bar{a}b}(s)+2s^{c\bar{d}}\hat{K}_{\bar{a}c}(s)\hat{K}_{\bar{d}b}(s)$ (28)
describing the metric of the clock subspace $\mathscr{H}_{T}$ evolves with respect to the one-dimensional object’s coordinate forming $\mathscr{H}_{X}$, in which $\hat{R}_{\bar{a}b}(s),\hat{K}_{\bar{a}b}(s)$ and $\hat{K}(s)$ are the intrinsic Ricci curvature, extrinsic curvature and scalar extrinsic curvature of $\mathscr{H}_{T}$, respectively.
Let’s examine the physical interpretation of the metric solutions of the equations. During the instrument calibration prior to measurement, mutually entangled quantum subsystems under-study $|X_{i}\rangle\in\mathscr{H}_{X}$ and the quantum instrument $|T_{a}\rangle\in\mathscr{H}_{T}$ are prepared. The entangled state they form can be written as 12. During the measurement process following instrument calibration, there is no further interaction between the quantum subsystem under-study and the quantum instrument. The coefficients $C_{i}$ and initial conditions for $h_{\bar{i}j}$ are established during the instrument calibration process and do not change with respect to any (hypothetical) external parameters afterward. The metric $h_{\bar{i}j}$ merely reflects the relative changes between the quantum subsystem under-study $X$ and the quantum instrument $T$. When the quantum instrument is in a certain quantum state $|T\rangle$ , the solutions to the equation provide us with the metric $h_{\bar{i}j}$ or basis vectors $|X_{i}\rangle$ of the subsystem under-study that is entangled with its basis vectors $|T_{a}\rangle$. Conversely, when the system is in a certain quantum eigenstate $|X\rangle$, the equation (28) yields the metric $s_{\bar{a}b}$ or basis vectors $|T_{a}\rangle$ that is entangled with its basis vectors $|X_{i}\rangle$. Thus, by utilizing the metric $s_{\bar{a}b}$ of eq.(28) or basis vectors $|T_{a}\rangle$, we can perform corresponding conditional projections on the entangled state $|X,T\rangle$ vector
$\langle T|X,T\rangle=\sum_{i}C_{i}\langle T|T_{i}\rangle|X_{i}\rangle=\sum_{i}C_{i}t^{i*}|X_{i}\rangle=\sum_{i}\frac{C_{i}}{t_{i}^{*}}|X_{i}\rangle,\quad\left(\textrm{repeated}\>i\>\textrm{not}\>\textrm{summate}\right)$ (29)
in which the expressions $|X,T\rangle=\sum_{i}C_{i}|X_{i}\rangle|T_{i}\rangle$ and the conditional projection state $|T\rangle=\sum_{a}t^{a}|T_{a}\rangle$ are used. The coefficient $\frac{C_{i}}{t_{i}^{*}}$ preceding $|X_{i}\rangle$ represents the relative amplitude for the quantum subsystem under-study in the state $|X_{i}\rangle$ when the quantum instrument is in a certain conditional quantum state $|T\rangle$. Since $C_{i}$ is determined by the initial conditions of instrument calibration and remains fixed after the preparation of the entangled state, the evolution equation (27) describes the evolution of the basis vectors $|X_{i}\rangle$ with respect to $|T_{a}\rangle$. That is, the solutions to the equation provide the correspondence between $|X_{i}\rangle$ and $t_{i}^{*}$, thereby establishing the relationship between the relative amplitude $\frac{C_{i}}{t_{i}^{*}}$ and $t_{i}^{*}$.
Thus, in this theory, there is no concept of state vector evolution as found in standard quantum mechanics (i.e., the unitary rotation of state vectors with respect to external parameters). Here, the entangled state vector $|X,T\rangle$ does not evolve (w.r.t. external parameter), instead, it is the basis vectors $|X_{i}\rangle$ within the Hilbert subspace that evolve relative to another subspace $|T_{a}\rangle$, causing the projection $\langle T|X,T\rangle$ to appear as if $|X_{i}\rangle$ is evolving. Precisely because only the basis vectors undergo relative evolution, this evolution is, in fact, universal or purely geometric. We note that the evolution equation (27) or (28) does not depend on the system’s mass nor explicitly contain the Planck constant (just as the Berry phase does not explicitly involve the Planck constant). The mass dependence and the Planck constant emerge when the equation (27) or (28) recovers the dynamic Schrödinger equation through the separation of fast-changing and slow-changing parts in the basis vectors $|X_{i}\rangle$ (see next section). In this sense, the relative changes between basis vectors on different subspaces described by the evolution equation are purely geometric Berry-type phase changes, rather than being divided into a dynamical part and a geometric (Berry-type) part with respect to external parameter as the standard quantum mechanics has shown.
By employing the metric solution $s_{\bar{a}b}$ of eq.(28) to perform a partial trace of the density matrix $\rho_{X,T}=|X,T\rangle\langle X,T|$ over the $\mathscr{H}_{T}$ component (since the entangled state $|X,T\rangle$ is not in Schmidt decomposition form, the partial trace generally requires contraction using the subspace metric rather than simply summing the diagonal elements of the subspace), that is, by contracting the density matrix with the metric of the subspace $\mathscr{H}_{T}$, we obtain
$\textrm{Tr}_{T}\left(\rho_{X,T}\right)=\langle T|X,T\rangle\langle X,T|T\rangle=\sum_{i,j}s_{\bar{i}j}C_{i}C_{j}^{*}|X_{i}\rangle\langle X_{j}|=\sum_{i}\frac{C_{i}C_{i}^{*}}{t_{i}^{*}t_{i}}|X_{i}\rangle\langle X_{i}|=\sum_{i}\left|\frac{C_{i}}{t_{i}}\right|^{2}|X_{i}\rangle\langle X_{i}|$ (30)
The diagonal elements of the matrix obtained after performing the partial trace directly represent the relative probabilities (or conditional probabilities) for the under-study subsystem in the state $|X_{i}\rangle$ when the quantum instrument is in the quantum state $|T\rangle$, yielding $\textrm{P}\left(X_{i}|T\right)=\left|\frac{C_{i}}{t_{i}}\right|^{2}$. As stressed before, the relative probability $\left|\frac{C_{i}}{t_{i}}\right|^{2}$ is not necessarily the standard absolute probability $|x_{i}|^{2}=h_{\bar{i}i}(t)$ being a solution of eq.(27). Only in the separable limit of the entangled state, it recovers the standard absolute probability $\textrm{P}\left(X_{i}|T\right)\rightarrow\textrm{P}\left(X_{i}\right)=|x_{i}|^{2}=h_{\bar{i}i}(t)$.
With the aid of this relative probability or the density matrix, the quantum expectation value of an Hermitian mechanical operator $\mathbf{O}(X)$, defined in $\mathscr{H}_{X}$ subspace, measured when the measuring instrument is in the quantum state $|T\rangle$, can now be expressed as
$\textrm{Tr}_{T}\left[\rho_{X,T}\mathbf{O}(X)\right]=\sum_{i}\left|\frac{C_{i}}{t_{i}}\right|^{2}\langle X_{i}|\mathbf{O}|X_{i}\rangle$
Similarly, the converse can also be obtained. When the under-study subsystem is in the quantum state $|X\rangle$, the relative probability of the quantum instrument being in the state $|T_{i}\rangle$ is $\textrm{P}\left(T_{i}|X\right)=\left|\frac{C_{i}}{x_{i}}\right|^{2}$. Correspondingly, the quantum expectation value of the mechanical operator $\mathbf{O}(T)$, defined in $\mathscr{H}_{T}$, under this relative probability is given by $\textrm{Tr}_{X}\left[\rho_{X,T}\mathbf{O}(T)\right]=\sum_{i}\left|\frac{C_{i}}{x_{i}}\right|^{2}\langle T_{i}|\mathbf{O}|T_{i}\rangle$.
We observe that the relative probabilities naturally emerge during the process of taking partial traces over subsystems for the entangled state $|X,T\rangle$. This differs from the partial tracing of entangled states in Schmidt decomposition as presented in standard quantum mechanics. Firstly, for entangled states in standard Schmidt decomposition, partial tracing by the Kronecker’s delta, the two reduced density matrices obtained by taking partial traces over two distinct subsystems share the same non-zero eigenvalues. In contrast, in our case, taking partial traces over different subsystems by each non-trivial metric yields different relative probabilities. Secondly, the subsystem basis vectors of the entangled state in the Schmidt decomposition are orthonormal, whereas the subsystem basis vectors of our prepared entangled states are not orthonormal but instead define a non-trivial subspace metric. This also leads to the fact that the metric of our subsystems satisfies the general evolution equation 27 or 28 for local basis vectors, rather than being associated with the constant-curvature Fubini-Study metric (which satisfies the constant-curvature Einstein equation for complex projective spaces), as is the case for Schmidt’s entangled states.
V Recovering the Schrödinger Equation
The evolution equations (27) or (28) plays a role analogous to that of the standard Schrödinger equation in describing the evolution of quantum states with respect to absolute external parameters. Therefore, in the following, we will start from the equation (27) and make some approximations to return to the Schrödinger equation limit. First, we linearize the Ricci curvature as $R_{\bar{i}j}\approx-\frac{1}{2}\Delta_{x}h_{\bar{i}j}$, where $\Delta_{x}$ is a linearized the Laplacian-Beltrami operator with respect to $x_{i}$, then eq.(27) becomes
$\frac{\partial^{2}h_{\bar{i}j}}{\partial t^{2}}=\Delta_{x}h_{\bar{i}j}+h^{k\bar{l}}\frac{\partial h_{\bar{i}k}}{\partial t^{a}}\frac{\partial h_{j\bar{l}}}{\partial t_{a}}-K_{a}\frac{\partial h_{\bar{i}j}}{\partial t_{a}}$ (31)
Using $h_{\bar{i}j}=\langle X_{i}|X_{j}\rangle$, and by equating the corresponding terms on both sides of the equation, it can be transformed into an equation in terms of the ket vector $|X_{i}\rangle$
$\frac{\partial^{2}|X_{i}\rangle}{\partial t^{2}}+K_{a}\frac{\partial|X_{i}\rangle}{\partial t_{a}}=\Delta_{x}|X_{i}\rangle+h^{k\bar{l}}\frac{\partial h_{\bar{i}k}}{\partial t^{a}}\frac{\partial|X_{l}\rangle}{\partial t_{a}}$ (32)
If the ket vector $|X_{i}\rangle$ now has a fast-changing characteristic frequency $\omega^{a}$ with respect to $t_{a}$, we will split $|X_{i}\rangle$ into two parts: a fast-changing part and a slow-changing part, i.e. $|X_{i}\rangle=e^{-i\omega^{a}t_{a}}|e_{i}\rangle$, where $|e_{i}\rangle$ represents the slow-changing part compared to $\omega^{a}$, such that $i\frac{\partial|e_{i}\rangle}{\partial t^{a}}\ll\omega_{a}|e_{i}\rangle$. Then we obtain
$\frac{\partial|X_{i}\rangle}{\partial t_{a}}=e^{-i\omega_{a}t^{a}}\left(\frac{\partial}{\partial t_{a}}-i\omega^{a}\right)|e_{i}\rangle$ (33)
If we consider the scalar extrinsic curvature $K_{a}=\frac{1}{2}h^{\bar{i}j}\frac{\partial h_{\bar{i}j}}{\partial t^{a}}$ changes also very slowly
$\frac{\partial}{\partial t^{b}}K_{a}\approx 0,\quad\frac{\partial}{\partial t^{b*}}K_{a}\approx 0$ (34)
Since $K_{a}=\frac{\partial}{\partial t^{a}}\ln\sqrt{\det h_{\bar{i}j}}$, the slow-changing approximation of this scalar extrinsic curvature is equivalent to the local volume of the subspace, $\sqrt{\det h_{\bar{i}j}}$, being approximately constant. This is a necessary condition for reverting to the unitarity of the standard quantum mechanics. It is also equivalent to the condition for returning to an inertial frame where no inertial forces are present (see next section).
Now the left hand side of the equation becomes
$\frac{\partial^{2}|X_{i}\rangle}{\partial t^{2}}+K_{a}\frac{\partial|X_{i}\rangle}{\partial t_{a}}\approx e^{-i\omega_{a}t}\left(\frac{\partial}{\partial t_{a}}-i\omega^{a}\right)\left(\frac{\partial}{\partial t^{a}}-i\omega_{a}+K_{a}\right)|e_{i}\rangle$ (35)
in which $\omega_{a}\omega^{a}=\omega^{2}$. When $|e_{i}\rangle$ is slow-changing, i.e. $\frac{\partial^{2}|e_{i}\rangle}{\partial t^{2}}\ll-2i\omega_{a}\frac{\partial|e_{i}\rangle}{\partial t_{a}}$, the term $\frac{\partial^{2}}{\partial t^{2}}|e_{i}\rangle$ can be neglected, the left hand side of the equation is approximately given by
$\frac{\partial^{2}|X_{i}\rangle}{\partial t^{2}}+K_{a}\frac{\partial|X_{i}\rangle}{\partial t_{a}}\approx e^{-i\omega_{a}t^{a}}\left[-2i\omega^{a}\left(\frac{\partial}{\partial t^{a}}+\frac{1}{2}K_{a}\right)-\omega^{2}\right]|e_{i}\rangle\equiv e^{-i\omega_{a}t^{a}}\left[-2i\omega^{a}\frac{D}{Dt^{a}}-\omega^{2}\right]|e_{i}\rangle$ (36)
Similarly, an independent equation satisfied by the bra vector $\langle X_{i}|$ can also be derived from the metric equation (32). Therefore, the bra vector $\langle X_{i}|$ and the ket vector $|X_{i}\rangle$ are mutually independent state vectors. Assuming that $\langle X_{i}|$ varies slowly compared to $|X_{i}\rangle$$\frac{\partial}{\partial t_{a}}\langle X_{i}|\ll-i\omega_{a}\langle X_{i}|$, we have
$\frac{\partial h_{\bar{i}k}}{\partial t_{a}}=\frac{\partial}{\partial t_{a}}\langle X_{i}|X_{k}\rangle=\frac{\partial\langle X_{i}|}{\partial t_{a}}|X_{k}\rangle+\langle X_{i}|\frac{\partial|X_{k}\rangle}{\partial t_{a}}\approx\langle X_{i}|\frac{\partial|X_{k}\rangle}{\partial t_{a}}=e^{-i\omega_{a}t^{a}}\langle X_{i}|\left(\frac{\partial}{\partial t^{a}}-i\omega_{a}\right)|e_{k}\rangle$
Since the slow-changing part of the ket vector satisfies $\frac{\partial}{\partial t_{a}}|e_{k}\rangle\ll-i\omega^{a}|e_{k}\rangle$, we neglect the $\frac{\partial}{\partial t_{a}}|e_{k}\rangle$ term, yielding
$\frac{\partial h_{\bar{i}k}}{\partial t_{a}}\approx\left(-i\omega^{a}\right)h_{\bar{i}k}$
That is, the variation in the metric $h_{\bar{i}k}$ is approximately dominated by the fast-changing part of the ket vector $|X_{k}\rangle$, while $\langle X_{i}|$ serves as a slow-changing basis, which is distinct from $\langle X^{i}|$, which is the fast-changing basis dual to $|X_{i}\rangle$. This can also be seen from the fact that, by $\langle X^{i}|X_{j}\rangle=\delta_{j}^{i}$,
$\frac{\partial\langle X^{i}|}{\partial t_{a}}|X_{j}\rangle=-\langle X^{i}|\frac{\partial|X_{j}\rangle}{\partial t_{a}}\approx i\omega_{a}\delta_{j}^{i}$ (37)
so $\langle X^{i}|$ is fast-changing. From alternative point of view, since $\langle X^{i}|=h^{\bar{i}j}\langle X_{j}|$ and $\langle X_{j}|$ is now a slow-changing basis, the fast-changing property of $\langle X^{i}|$ is entirely attributed to the fact that the metric $h^{\bar{i}j}$ is fast-changing, specifically $\frac{\partial h^{\bar{i}j}}{\partial t_{a}}\approx\left(i\omega^{a}\right)h^{\bar{i}j}$.
Based on the above results, the right-hand side of equation (32) now approximately yields
$\Delta_{x}|X_{i}\rangle+h^{k\bar{l}}\frac{\partial h_{\bar{i}k}}{\partial t^{a}}\frac{\partial|X_{l}\rangle}{\partial t_{a}}\approx e^{-i\omega_{a}t^{a}}\Delta_{x}|e_{i}\rangle+h^{k\bar{l}}\left(-i\omega^{a}\right)h_{\bar{i}k}e^{-i\omega_{a}t^{a}}\left(-i\omega_{a}\right)|e_{l}\rangle=e^{-i\omega_{a}t^{a}}\left(\Delta_{x}-\omega^{2}\right)|e_{i}\rangle$ (38)
Finally, equating the left hand side (36) and right hand side (38), the eq.(32) can be approximately given by a Schrödinger-like equation
$i\frac{D|e_{i}\rangle}{Dt^{a}}\approx-\frac{1}{2\omega^{a}}\Delta_{x}|e_{i}\rangle$ (39)
in which
$\frac{D}{Dt^{a}}\equiv\frac{\partial}{\partial t^{a}}+\frac{1}{2}K_{a}$ (40)
is a covariant derivative, in which the scalar extrinsic curvature $K_{a}$ now plays the role of a connection for the parallel transport of the slow-changing basis $|e_{i}\rangle$ on the Hilbert space $\mathscr{H}_{T}$. If we regard the complex coordinates $t_{a}$ and $x_{i}$ in this equation as real coordinates (mathematically, it is always possible to appropriately choose coordinates or basis vectors $|T_{a}\rangle$ and $|X_{i}\rangle$ in the local quantum state subspace to render the local coordinates $t_{a}$ and $x_{i}$ real), then, in terms of the real coordinates, this equation formally becomes a Schrödinger equation.
To sum up, we have demonstrated that the equation (27) or its linearized approximations (31) or (32) describe the evolution of the metric $h_{\bar{i}j}$ or the basis vectors $|X_{i}\rangle$ in the subspace $\mathscr{H}_{X}$ with respect to the entangled state vector $|T\rangle$ in $\mathscr{H}_{T}$. This is because we observe that the slow-changing part $|e_{i}\rangle$ of the basis vectors $|X_{i}\rangle$ satisfies an approximate “free” Schrödinger evolution equation, where the fast-changing characteristic frequency $\omega^{a}$ a plays a role analogous to the mass $M=\hbar\omega^{a}$, in which the mass and the Planck’s constant emerge, and the potential term in the Schrödinger equation is embodied in the connection (scalar extrinsic curvature $K_{a}$) of the covariant derivative.
VI Non-Inertial Effects
Because in this quantum theory, the evolution of quantum states is no longer relative to the external Newtonian time of an absolute inertial frame, but rather to the physical clock time in a general (non-inertial) quantum reference system, and we have also observed that this “free” Schrödinger evolution equation actually contains a potential term, which generates an additional “external force”, so this quantum theory will automatically incorporate some effects of non-inertial frames. As a result, certain “inertial forces” will automatically emerge from the covariant derivative and the connection $K_{a}$. This is analogous to the classical case that the Newton’s law of motion in inertial frame is generalized to the geodesic equation in general non-inertial frames, and the inertial force (and gravity force) felt by an object is encoded in the connection of the geodesic equation.
Through this connection $K_{a}$ in (40), if the state vector $|X\rangle$ moves along a path $c$ on $t\in\mathscr{H}_{T}$, it acquires an additional non-integrable phase factor $e^{\varTheta}$, (in general $\varTheta\in\mathbb{C}$)
$|X^{\prime}\rangle=e^{\varTheta}|X\rangle$ (41)
where
$\varTheta=\frac{1}{2}\int_{c}dt^{a}K_{a}$ (42)
and $\varTheta$ not only encompasses the dynamics phase change (coming from the dynamic Schrödinger-like equation) of the slow-changing basis $|e_{i}\rangle$ but also includes other additional phase changes of the state $|X\rangle$. If the inner product of states is regarded as a certain kind of distance, then factor $e^{\varTheta}\in\mathbb{C}$ can also be interpreted as the finite (complex) geodesic distance between the quantum states $|X\rangle$ and $|X^{\prime}\rangle$. When $c$ is a closed loop, i.e. when the state returns to its starting point, the phase change contributed by the fast-changing dynamical part in the phase disappears, leaving behind a phase change related to the intrinsic geometry of the closed loop $c$. Since this closed loop lies within the Hilbert subspace c rather than in a parameter space, in this sense, the geometric phase is more akin to the Aharonov-Anandan phase (a generalization of the Berry phase) induced by a closed loop in the projected quantum state space. The phase $\varTheta$ and the path accumulation of the potential term (the connection or extrinsic curvature $K_{a}$) thus generally constitute a non-integrable phase. The derivative of the “potential” $K_{a}$ gives rise to a certain kind of “external force”
$F_{\bar{a}b}\equiv\frac{\partial K_{b}}{\partial t^{a*}}$ (43)
where $F_{\bar{a}b}$ reflects a kind of "inertial force" that arises due to the choice of quantum reference system when $|X\rangle$ is considered relative to $|T\rangle$. Under the approximation (34), we return to the situation of an approximately inertial frame where it is inertial-force-free.
Furthermore, we can define another metric
$g_{\bar{a}b}\equiv\left\langle\frac{\partial X^{i}}{\partial t^{a}}\right|\left.\frac{\partial X_{i}}{\partial t^{b}}\right\rangle\equiv G_{\bar{a}b}+\frac{i}{2}\Omega_{\bar{a}b}$ (44)
in which $G_{\bar{a}b}$ and $\Omega_{\bar{a}b}$ are the real and imaginary part of the metric. As a metric coming from the comparison between $|X^{i}\rangle$ and $|T^{a}\rangle$, it can be interconverted with the previous metrics $s_{\bar{a}b}$ and $h_{\bar{i}j}$ through a tensor coordinate transformation, for example,
$g_{\bar{a}b}=\left\langle\frac{\partial X^{i}}{\partial t^{a}}\right|\left.\frac{\partial X^{j}}{\partial t^{b}}\right\rangle h_{i\bar{j}}=\left\langle\frac{\partial T_{a}}{\partial x_{i}}\right|\left.\frac{\partial T_{b}}{\partial x_{j}}\right\rangle h_{i\bar{j}}=\left\langle\frac{\partial T_{a}}{\partial t_{c}}\right|\left.\frac{\partial T_{b}}{\partial t_{d}}\right\rangle s_{\bar{c}d}$ (45)
In contrast to the previous interpretation of $s_{\bar{a}b}$ and $h_{\bar{i}j}$ as intrinsic metrics of the subspaces, $g_{\bar{a}b}$ can also be interpreted as a relative metric between $\mathscr{H}_{T}$ and $\mathscr{H}_{X}$, in certain sense, representing the pull-back of the full-space metric $q_{\bar{I}J}$ onto the subspaces.
It can be shown that this metric $g_{\bar{a}b}$ is, in fact, equivalent to the “inertial force”
$F_{\bar{a}b}=g_{\bar{a}b}$ (46)
From (41), it can be observed that the real part of the metric, $G_{\bar{a}b}=G_{b\bar{a}}=\textrm{Re}\left(g_{\bar{a}b}\right)$, provides the (real) distance between two infinitesimally separated state vectors $|X(t+dt)\rangle$ and $|X(t)\rangle$. Since the state vector $|X\rangle$, as a local fiber in the non-trivial fiber bundle, can only be normalized locally at $t$, both the length and the phase of the state vector $|X\rangle$ will change relative to $|T\rangle$ as $t$ varies. Since
$\langle X^{i}(t+dt)|X_{i}(t)\rangle=\langle X^{i}(t)|X_{i}(t)\rangle+\frac{1}{2}\left\langle\frac{\partial X^{i}}{\partial t^{a}}\right|\left.\frac{\partial X_{i}}{\partial t^{b}}\right\rangle dt^{a*}dt^{b}+...=e^{d\varTheta}=1+d\varTheta+$... (47)
so the real part $G_{\bar{a}b}$ corresponds to
$\textrm{Re}\left(d\varTheta\right)$ $=\frac{1}{2}\textrm{Re}\left\langle\frac{\partial X^{i}}{\partial t^{a}}\right|\left.\frac{\partial X_{i}}{\partial t^{b}}\right\rangle dt^{a*}dt^{b}=\frac{1}{2}G_{\bar{a}b}dt^{a*}dt^{b}$
$=\left\|\langle X^{i}(t)|X_{i}(t)\rangle-\langle X^{i}(t+dt)|X_{i}(t)\rangle\right\|=\left\|1-\langle X^{i}(t+dt)|X_{i}(t)\rangle\right\|\in[0,1]$ (48)
Comparing the above equation and (42), we see the relation
$\textrm{Re}\left(d\varTheta\right)=\frac{1}{2}G_{\bar{a}b}dt^{a*}dt^{b}=\textrm{Re}\left(\frac{1}{2}K_{b}dt^{b}\right)$ (49)
so we can conclude that, at certain approximation, the real part of the metric is nothing but the real part of the “inertial force”
$G_{\bar{a}b}=\textrm{Re}\left(\frac{\partial K_{b}}{\partial t^{a*}}\right)=\textrm{Re}\left(F_{\bar{a}b}\right)$ (50)
Furthermore, if the real and imaginary parts of the metric are interconnected by the complex structure operator $J$ of the state space (which satisfies the eigenvalue equations $J\frac{\partial}{\partial t^{a}}=i\frac{\partial}{\partial t^{a}}$ and $J\frac{\partial}{\partial t^{a*}}=-i\frac{\partial}{\partial t^{a*}}$), i.e.
$JG_{\bar{a}b}=-\Omega_{\bar{a}b},\quad J\Omega_{\bar{a}b}=G_{\bar{a}b}$ (51)
so the imaginary part is given by
$\Omega_{\bar{a}b}=2\textrm{Im}\left(\frac{\partial K_{b}}{\partial t^{a*}}\right)=2\textrm{Im}\left(F_{\bar{a}b}\right)$ (52)
The imaginary part of the metric is also the imaginary part of the “inertial force”. Therefore we have the relation eq.(46).
We observe that when the phase $\varTheta$ is integrable, or equivalent speaking, the state space is Kähler, there exists a locally defined Berry curvature $B_{\bar{a}b}$, which is defined as the exterior derivative of the connection $K_{a}$, i.e. $B=dK$
$B_{\bar{a}b}=\frac{\partial K_{b}}{\partial t^{a*}}-\frac{\partial K_{a}}{\partial t^{b*}}=F_{\bar{a}b}-F_{a\bar{b}}=F_{\bar{a}b}-F_{\bar{a}b}^{*}$ (53)
Comparing this equation with the definition $\Omega_{\bar{a}b}\equiv g_{\bar{a}b}-g_{\bar{a}b}^{*}$ by using (46), we conclude that $\Omega_{\bar{a}b}$ can be interpreted as the Berry curvature (or field strength) $B_{\bar{a}b}$
$\Omega_{\bar{a}b}\doteq B_{\bar{a}b}$ (54)
The symbol $\doteq$” means that the equality holds only when the Kähler form of the state space is closed, in other words, the state space is strictly a Kähler manifolds or the phase $\varTheta$ is integrable (see below). However, in general, the geometry is only almost Kähler but not strictly so.
It is worth stressing that only when the phase $\varTheta$ is integrable, (e.g. in a source-free case of the electromagnetic fields where the field strength is closed $dB=0$), $d\Omega=0$, where $\Omega=\frac{i}{2}B_{\bar{a}b}dt^{a*}\wedge dt^{b}$ as the Kähler form of the state space is closed, in this case, the state space is strictly Kähler.
Next, let’s discuss some physical implications of the real and imaginary parts of the non-integrable phase $\varTheta$. We observe that $\textrm{Re}\left(d\varTheta\right)=\left\|1-\langle X^{i}(t+dt)|X_{i}(t)\rangle\right\|=\frac{1}{2}G_{\bar{a}b}dt^{a*}dt^{b}$ can also be interpreted as an infinitesimal stretching of the length of the state vector $|X\rangle$ (i.e. $\left\||X\rangle\right\|$), which makes the evolution of $|X\rangle$ with respect to $|T\rangle$ no longer unitary in general. Consequently, previously we could only derive the Schrödinger equation with unitary evolution under the assumption of slow-changing of the scalar extrinsic curvature (34) $G_{\bar{a}b}=\textrm{Re}\left(\frac{\partial K_{b}}{\partial t^{a*}}\right)\approx 0$. This is also an inevitable consequence of the relative evolution of subsystems in entangled states, resulting from the quantum superposition (quantum second-moment spreading (Luo2015Dark, ; Luo:2023eqf, ; 2024arXiv240809630L, )) of the $T$ subsystem that serves as the reference frame.
This breakdown of unitarity can also be interpreted as an effect of additional second-order fluctuations or extra broadening. Considering the second-moment spreading of the effective energy $E$ corresponding to the second-moment fluctuations of $t$, namely $\langle\delta E^{2}\rangle=\langle E^{2}\rangle-\langle E\rangle^{2}$, it leads to $\langle X^{i}(t+dt)|X_{i}(t)\rangle=1-\frac{1}{2}\langle\delta E^{2}\rangle dt^{2}$. Therefore, the metric $g_{\bar{a}b}$ is directly determined by the second-moment quantum fluctuations of $t$ or the second-moment fluctuations of the effective energy, i.e.
$g_{\bar{a}b}dt^{a*}dt^{b}\sim\langle\delta E^{2}\rangle|dt|^{2}$ (55)
This is nothing but the Anandan-Aharonov theorem (Anandan:1990fq, ). Hence, this “inertial force” $F_{\bar{a}b}=g_{\bar{a}b}\sim\langle\delta E^{2}\rangle$ can also be regarded as arising from the quantum fluctuations of energy or quantum fluctuations of the clock reference frame. We also observe that the breakdown of unitarity caused by the second-order fluctuations of the quantum clock and the broadening of spectral lines (Luo:2023eqf, ) both are coming from the additional “inertial force” effect. Conversely, the effect of the “inertial force” always seems to be associated with changes in the norm of the state vector or the breakdown of quantum unitarity. This non-unitarity resulting from the broadening of the quantum clock or spectral lines will be generalized to more general quantum spacetime reference frames in (Luo:2015pca, ; Luo:2019iby, ; Luo:2021zpi, ; 2024arXiv240809630L, ), and is the fundamental reason of coordinate transformation anomalies, the cosmological constant, and gravity in general quantum spacetime reference frames.
Similarly, since the imaginary part of the phase $\varTheta$ and metric $g_{\bar{a}b}$, $\Omega_{\bar{a}b}$ is related to the Berry curvature,
$d\theta=\textrm{Im}\left(d\varTheta\right)=\frac{1}{4}\Omega_{\bar{a}b}dt^{a*}\wedge dt^{b}\doteq\frac{1}{4}B_{\bar{a}b}dt^{a*}\wedge dt^{b}$ (56)
so the imaginary part describes an infinitesimal phase change of $|X\rangle$, different from that the real part gives rise to the length change of $|X\rangle$ (48) and (49). The integration along the closed loop $c$ is given by
$\theta[c]=\textrm{Im}\left(\varTheta\right)=\frac{1}{2}\textrm{Im}\oint_{c}dt^{a}K_{a}\doteq\frac{1}{4}\iint_{s}B_{\bar{a}b}dt^{a*}\wedge dt^{b}$ (57)
It describes the geometric phase change of the state vector $|X\rangle$ as it traverses a closed loop $c=\partial s$ on $\mathscr{H}_{T}$.
To sum up, we see that the “inertial force” $F_{\bar{a}b}$ coming from the potential $K_{a}$ in eq.(39) is divided into two parts: the real part can be seen as providing a “stretching force” $G_{\bar{a}b}$ related to the change in the length of the state vector, while the imaginary part gives rise to a “gauge force” $B_{\bar{a}b}$ related to phase changes (we know that gauge fields only alter the phase of the state vector without changing its magnitude). The “stretching force” modifies the length of the state vector, while the “gauge force” modifies its phase. In this sense, eq.(39) in general coordinate system automatically introduces a “stretching force” and a “gauge force” at the quantum level. The profound relationship between the imaginary part and the gauge force is also quite intriguing, and can be found in other books discussing the geometric phase in conventional quantum theory, as well as in the author’s previous articles. Our previous primary interest lies in the “stretching force”, because it actually manifests as an Einstein-gravity in the quantum reference frame case (Luo:2021zpi, ; 2024arXiv240809630L, ) being the generalization of the quantum clock. Generally speaking, as long as a general complex connection (with both real and imaginary parts) is introduced in the general transformation of the quantum state relative to the quantum clock, resulting in a non-integrable factor $e^{\varTheta}$ for the quantum state, it inevitably introduces both gravity and gauge force through such a general coordinate transformations. Whether this achieves some form of unification between gravity and gauge force is a more profound question that lies beyond the scope of this paper.
As observed earlier, the real part $G_{\bar{a}b}$ describes the second-order fluctuations in the length of vectors on $\mathscr{H}_{X}$ relative to those on $\mathscr{H}_{T}$, reflecting the degree of unitarity violation associated with changes in the norm of the state vector. If we note that the scalar extrinsic curvature actually describes the first-order variation of the local volume element of $\mathscr{H}_{X}$ relative to the vector $|T\rangle$,
$K_{b}=\frac{\partial}{\partial t^{b}}\ln\sqrt{\det h_{\bar{i}j}}$ (58)
so the real part of the metric can be given by
$G_{\bar{a}b}=\textrm{Re}\left(\frac{\partial K_{b}}{\partial t^{a*}}\right)=\textrm{Re}\left(\frac{\partial^{2}}{\partial t^{a*}\partial t^{b}}\ln\sqrt{\det h_{\bar{i}j}}\right)$ (59)
It naturally follows that the (logarithmic) magnitude of the local volume element of $\mathscr{H}_{X}$ undergoes second-order variations with respect to $|T\rangle$. The changes in the local volume of $\mathscr{H}_{X}$ are precisely caused by the quadratic correction to the length of its quantum state due to second-order fluctuations. Under the slow-changing approximation of the extrinsic curvature (34), the local volume is approximately conserved, and unitarity approximately holds. Meanwhile, the imaginary part $\Omega_{\bar{a}b}$, acting as the Berry curvature, describes the second-order phase change of the state vector on $\mathscr{H}_{X}$ relative to that on $\mathscr{H}_{T}$.
We observe from (59) that the logarithm of the local volume element of $\mathscr{H}_{X}$, namely $\ln\sqrt{\det h_{\bar{i}j}}$, formally plays a role analogous to the Kähler potential, meaning that the complex metric can be expressed in terms of a single volume form
$g_{\bar{a}b}=\partial_{\bar{a}}\partial_{b}\ln\sqrt{\det h_{\bar{i}j}}$ (60)
If the phase $\varTheta=\frac{1}{2}\int_{c}dt^{a}K_{a}$ is integrable (independent to the path $c$), meaning that the local Berry curvature $B=B_{\bar{a}b}dt^{a*}\wedge dt^{b}$ is closed ($dB=d^{2}K=0$), and so if this local Berry curvature corresponds to the imaginary part $\Omega_{\bar{a}b}$ of the local metric, the quantum state space equipped with this Hermitian complex metric must be a Kähler space. In this case, the gauge field strength 2-form $\frac{i}{2}B_{\bar{a}b}dt^{a*}\wedge dt^{b}\doteq\frac{i}{2}\Omega_{\bar{a}b}dt^{a*}\wedge dt^{b}$ serves as the Kähler form, reflecting the symplectic structure of the Kähler space. A Kähler space is one that simultaneously possesses a Riemannian structure, a symplectic structure, and is compatible with its complex structure. When the quantum state space is a Kähler space, or equivalently, the phase $\varTheta$ is integrable, due to the enhanced symmetry of the Kähler metric, which can be represented by a single Kähler potential, many problems by solving the multi-component Ricci-flat Kähler-Einstein equations can be reduced to solving a single complex Monge-Ampère equation, which simplifies the problem of solving the Kähler-Einstein equations. When there exists a constant state space volume form $\sqrt{\det h_{\bar{i}j}}$, i.e. a constant Kähler potential $\ln\sqrt{\det h_{\bar{i}j}}$, this degenerates to the case of a unitary quantum mechanics, so the state space of the unitary standard quantum mechanics is Kähler, which is suggested by literature e.g. (Ashtekar:1997ud, ; Bengtsson:2006rfv, ). In particular, when the curvature of its subspace is proportional to the metric, $R_{\bar{a}b}=(N+1)g_{\bar{a}b}$, the quantum state space is homeomorphic to the maximally symmetric sphere $S^{2N+1}$ and this Kähler metric $g_{\bar{a}b}$ is known as the Fubini-Study metric (Pandya:2006mrv, ; Matassa:2020dnc, ). However, the subspace metric $h_{\bar{i}j}$ of interest to us generally does not possess constant curvature or a constant local volume form, so the subspace metric of the quantum state here is generally not Fubini-Study in this framework. Nonetheless, in general, our phase $\varTheta$ is not strictly integrable (depending on the path and history), and the Berry curvature can no longer be locally defined. In this case, the quantum state subspace of the framework is not strictly Kähler but can be regarded as an almost Kähler manifold.
VII Comparison with Some Past Works
Given the long-standing research on the relational formulations and internal evolution of quantum mechanics, coupled with the already vast body of literature on the subject, it has become impossible for us to fully grasp all the existing works. While our research may have been partially or indirectly inspired by some past studies in terms of major philosophical directions, the core ideas of our work are independently discovered. Our primary motivations for developing the theory of quantum reference frames are entirely different from those in previous literature. The initial few papers were motivated by the cosmological constant problem (Luo2014The, ; Luo2015Dark, ; Luo:2015pca, ) and quantum gravity, which must involve quantum spacetime as reference frame, whereas the present paper focuses solely on quantum mechanics which involves only quantum clock. Nevertheless, it was only later that we realized our thoughts converged with those of these past studies in certain aspects, albeit through different paths.
Our paper achieves results that are similar in many respects to those in the work by Page and Wootters (PhysRevD.27.2885, ). However, there are also numerous differences between our paper and their study. Firstly, in Page and Wootters’ article, the Hamiltonian still exists, and the evolution of operators is realized through it. In contrast, our formulation is purely covariant and no longer incorporates a Hamiltonian, since the presence of any Hamiltonian implicitly implies a preferred inertial frame which is a concept that the general relativity aims to discard from the outset. Therefore, the quantum reference system discussed in our article is a general one, where inertial frames hold no special status whatsoever. That is the reason why the general non-inertial effects appear which seem absence in Page-Wootters’ paper. We also observe that it is precisely the quantum fluctuations of the quantum reference system that dictate the non-existence of rigorous inertial frames at the quantum level. Secondly, although both our paper and theirs provide methods for calculating conditional probabilities, these methods are entirely distinct. In their work, the calculation of conditional probabilities is carried out by taking the partial trace via the projection of the quantum state of a specific quantum reference frame subspace which is flat. In other words, this quantum reference frame is still associated with some kind of flat inertial frame, rather than being a general quantum coordinate system. In our framework, conditional probability are computed via projections or partial traces with respect to a generally curved metric of the quantum reference frame subspace, featuring a more natural and covariant geometric construction. Moreover, the entangled state (12) prepared during calibration, is fixed, and does not evolve with respect to external Hamiltonian, instead, the relative evolution of the under-study $|X_{i}\rangle$ with respect to the clock subsystem arises solely from the projection metric $s_{\bar{a}b}\neq\delta_{\bar{a}b}$ of the subspace which only evolves internally. And through geometric method, the paper has found the dynamical equation governing the evolution of the general metric (28), which indirectly gives the time evolution of the conditional probability by the internal evolution of the subspace metric, completely different from Page-Wootters and related theory (PhysRevD.27.2885, ; 1986Quantum, ; Foti:2020erm, ; Woods2019pagewootters, ; qfns-48vq, ). Therefore, this article is not merely a simple re-formulation of the main results of the Page-Wootters framework, rather, it can be regarded as presenting a more general generalization, the new physics is those non-inertial effects coming from the general covariance without a Hamiltonian, just as that general relativity is not a simple re-formulation of Newton’s laws, the new physics is the gravity naturally incorporated.
Although our discussions only focus on the quantum clock (not a quantum spacetime), it is of cause, in essence, a quantum reference frame theory. However, frankly speaking, the results obtained in the existing literature on quantum reference frames (PhysRevD.30.368, ; Rovelli:1990pi, ; RevModPhys.79.555, ; Giacomini:2017zju, ; Loveridge:2017pcv, ; 2020Quant…4..225V, ; Hohn_2020, ; Giacomini:2020ahk, ) are still quite preliminary, in the sense that it is still a long way off from the real goal of this subject: quantum gravity. Compared our works with the existing literature of quantum reference frame, firstly, most current articles on quantum reference frames are still based on the description of quantum reference frames via Hamiltonian. As previously mentioned, this implicitly implies the selection of a preferred inertial frame, making it difficult for these theories to derive non-inertial effects, let alone naturally incorporate gravitational effects. On the other hand, current literature on quantum reference frames still lack a fundamental quantum equivalence principle as a guiding principle to re-interpret the universal quantum properties of material quantum reference frames as the universal properties of the abstract concept of quantum time itself or quantum spacetime itself. In our paper on the quantum equivalence principle (2024arXiv240809630L, ), we first point out that only non-dynamical universal part of quantum second-order moment fluctuation (not coming from its Hamiltonian or mass) of physical reference frames, can universally describe the second-order moments of quantum clocks or quantum spacetime, thereby reconciling the inherent contradiction between the standard Schrödinger equation of quantum mechanics and the equivalence principle (different from most of the literature on the subject, e.g. (Giacomini:2020ahk, ; Das:2023cfu, )). Therefore, our quantum reference frame theory is built upon a covariant theory with a zero Hamiltonian, which is different from most of the literature of the subject. Consequently, in this paper, such quantum clocks naturally give rise to non-inertial effects. Furthermore, we can also discuss here some comparisons between our quantum spacetime reference frame theory (e.g. (Luo:2021zpi, ; 2024arXiv240809630L, ; 2024chinarxiv, )) beyond the scope of quantum clock and the existing literature on the subject. In our quantum spacetime reference frame (a non-linear sigma model with zero Hamiltonian as a generalization of the quantum clock, which is more suitably treated by functional method rather than the canonical method), the diffeomorphism anomalies naturally lead to the emergence of gravity (Luo:2021zpi, ), and the cancellation of anomalies is connected to the cosmological constant—a feat that has not been truly achieved in the most literature on the quantum reference frame. Secondly, our quantum reference frame theory provides the renormalization group flow of the quantum spacetime (which is, to an approximation, a Ricci flow) and its associated monotonic entropy functional (Luo:2022statistics, ; Luo:2025cer, ), which gives an origin of gravitational entropy, and enabling a global control of a gravitational systems (since a covariant system like gravity lacks a global time coordinate and hence the concept of energy no longer provides a global control like usual dynamical system), a feat not achieved in the current quantum reference frame literature. Moreover, a general quantum coordinate transformation is given by a Jacobian in the functional integral approach which leads to diffeomorphism anomaly and gravity, or equivalently given by an integral transformation via a heat kernel of the conjugate heat flow equations which is coupled to the spacetime Ricci flow.
Of course, as a quantum reference frame theory, the philosophical interpretation of quantum mechanics presented in this paper is much closer to that of the relational quantum mechanics (Rovelli:1995fv, ; Rovelli:2021thu, ), and has little to do with the “relative state” in the so-called many-worlds interpretation, despite the semantic similarity between “relational” and “relative”—we do not wish to get bogged down in semantic quibbles here, and we also don’t want to get entangled in the philosophical disputes among various schools of thought derived from the relational interpretation. The fundamental viewpoint we advocate in this paper is that quantum states are relative (w.r.t. quantum apparatus), and physical properties can only be defined relatively. Hence, there is no global section (of the non-trivial fiber bundle) serving as a global wave function and moreover no global underlying “hidden variable”. The “relative state” is a local section on a non-trivial fiber bundle, whereas the many-worlds interpretation still posits a global wave function, with observers merely perceiving branches of this global wave function. Of course, unlike relational quantum mechanics, which merely offers a new interpretation of the existing mathematical formalism of quantum mechanics, we strive to more thoroughly incorporate the spirit of general relativity into the framework of quantum mechanics.
VIII Conclusions
We summarize the main conclusions of this paper as follows:
(1) This paper presents a quantum theory based on entangled or relative states rather than absolute quantum states. This quantum theory describes how a quantum system evolves relative to another quantum system, especially, describing a quantum state of the relationship between the quantum under-study system and that of the quantum measuring apparatus, without relying on any external absolute parameter. The relative interpretation serves as a physical interpretation of the entangled state, and the geometrical interpretation of the entangled state is in fact a non-trivial fiber bundle. The picture of the non-trivial fiber bundle interpretation of entangled state also constitutes an important motivation for our proposal that a locally curved state space geometric description is required to depict the relative evolution of quantum states.
(2) We gives a non-trivial fiber bundle interpretation to the entangled state $|X,T\rangle=\sum_{\tau}C_{\tau}|X(\tau)\rangle\otimes|T(\tau)\rangle$, in which $|T(\tau)\rangle$ is the local base space, $|X(\tau)\rangle$ is the fiber over the local base $|T(\tau)\rangle$, the entangled state $|X,T\rangle$ is then the fiber bundle, and the projection $T^{-1}:|X,T\rangle\rightarrow|X(\tau)\rangle$ from the fiber bundle to a section of the fiber gives rise to the relative amplitude of the under-study quantum state $|X(\tau)\rangle$ relative to the clock state $|T\rangle$. So there is a relative probability interpretation concerning the probability of state $|X(\tau)\rangle$ relative to the measuring apparatus state $|T\rangle$, rather than an absolute probability interpretation relative to external parameters. The relative or conditional probabilities derived from an entangled state reflect that our information about the entangled system is incomplete. When one subsystem of an entangled state learns, via a classical channel, the measurement result of the other subsystem, it can utilize this additional information to update our knowledge about this local subsystem in the sense of the Bayesian theorem. Only when the entangled state is separable or the measuring apparatus can be considered classical, then entangled state degenerates into a direct product state (the non-trivial fiber bundle degenerates to a trivial bundle), does this relative probability recovers the absolute probability of the standard quantum mechanics.
(3) The entangled state of the under-study state $|X\rangle$ and the clock apparatus state $|T\rangle$, prepared during calibration, does not evolve with respect to external parameters. However, the relative evolution of the under-study quantum state $|X\rangle$ relative to the clock state $|T\rangle$ in the intrinsic quantum theory can be described by a evolution equation of the total quantum state space coming from the Ricci-flat Kähler-Einstein equation. The evolution equation is a generally covariant equation that no longer refers to any external absolute parameters and constrained by a vanishing Hamiltonian, but only relies on internal relative descriptions. In this sense, this framework achieves a geometric formulation of a quantum theory describing the under-study quantum system and the quantum apparatus. This equation has a linear and non-relativistic approximation to the Schrödinger equation.
(4) From this framework, the evolution of quantum states is now the evolution of basis vectors of one subspace relative to those of another subspace. Although the entangled state (relative to external parameters) does not evolve, the projection components of the entangled state onto the basis vectors of the subspace evolve relative to the clock subspace. In this picture, there is no concept of a state vector evolving relative to unchanging basis vectors; instead, the entangled state vector remains unchanged while the basis vectors change. Consequently, there is no longer a distinction between quantum dynamical phases (where basis vectors remain unchanged and state vectors change) and geometric phases (where basis vectors change relative to external parameters). Now, the “dynamical phase” essentially also arises from the geometric phase due to changes in the basis vectors. In this sense, the quantum dynamical evolution is geometrized, as the Newton’s dynamical law is geometrized by the geodesic equation in general coordinate system.
(5) Within this framework, there is essentially no difference between the time evolution of the under-study system (by a quantum clock) and a measurement process of other aspects of the under-study system (by other quantum instruments). The distinction lies in the fact that the quantum clock provides a reference state with no interaction with the state of the under-study system, whereas other general measuring instruments offer other reference states with certain interaction with the state of the under-study system. Consequently, in certain sense, this framework offers a unified description that bridges the time evolution (which is described by the unitary Schrödinger equation in standard quantum mechanics) and the general quantum measurement process (which is depicted by a non-unitary collapse of wavefunction in the standard quantum mechanics).
(6) For the intrinsic quantum system of the measured quantum state $|X\rangle$ relative to the clock state $|T\rangle$, we observe the natural emergence of “inertial forces” associated with the choice of a general clock reference state. This “inertial force” $F_{\bar{a}b}$ is related to the non-trivial relative metric $g_{\bar{a}b}$ between the under-study quantum state $|X\rangle$ and the clock apparatus state $|T\rangle$. The “inertial force” and its quantum effects, derived from the general choice of reference frame, are issues of concern in (Luo:2021zpi, ). The real part of this “inertial force” brings about certain “stretching force” and second-order broadening of the entangled subsystem which leads to the destruction of quantum state unitarity and anomaly (volume changes in the subsystem quantum state space and information loss dependent on history), while the imaginary part brings about certain “gauge force” and geometric phases. When the phase factor is non-integrable, just as the geometric phase brought about by the imaginary part is history-dependent, the destruction of unitarity by the real part is also history- or path-dependent. This historical dependence of quantum non-unitary effects has already been observed in quantum field theory in curved spacetime. Throughout this paper, we will see that the quantum unitarity-violation effects and anomalies brought about by non-inertial frames are very common and general conclusions. However, the “inertial force” $F_{\bar{a}b}$ derived in this simple system does not yet exhibit gravitational characteristics. It is merely a general inertial force. The reason is that our current system is still a simple (single reference degree of freedom) quantum clock reference system. When extended to a multiple reference degrees of freedom, namely, a quantum spacetime reference system, the characteristics of an Einstein-gravity arising from the unitarity-violation and anomalies caused by this type of inertial force will become apparent (e.g (Luo:2021zpi, ; 2024arXiv240809630L, )), it automatically leading to a gravity theory coming from the general covariance of quantum reference systems. Our previous works on the functional integral approach to the quantum spacetime reference frame only focuses on the real part and its relation to the anomaly and gravity, while the effects of the imaginary part have not been taken into account, now it seems that the relative state approach also reveals the effects of the imaginary part and its relation to the Berry phase and “gauge force”.
(7) The relative evolution equation for the quantum state of the one-dimensional object’s position relative to a single-degree-of-freedom quantum clock reference system has already been a high dimensional evolution equation. Therefore, it can be imagined that if more degrees of freedom of the quantum spacetime reference system are involved, this canonical framework will become even more complicated. This canonical framework, based on differential equations for quantum state evolution, is more suitable for conceptually demonstrating its relationship with the standard quantum mechanical Schrödinger equation and its probability interpretation, helping us observe more effects that are not easily seen through the functional integral approach. On the other hand, the action and functional integral quantization framework is actually more amenable to practical mathematical calculations and to study quantum gravity. Each method has its own advantages and disadvantages. To facilitate comparison between these two methods, we will provide the functional method for a single-degree-of-freedom quantum clock in the Appendix for easy comparison. For the functional treatment of quantum spacetime reference systems with more degrees of freedom, please refer to (Luo:2021zpi, ; 2024arXiv240809630L, ; 2024chinarxiv, ).
Acknowledgements.
This work was supported in part by the National Science Foundation of China (NSFC) under Grant No.11205149, and the Scientific Research Foundation of Jiangsu University for Young Scholars under Grant No.15JDG153.
Appendix: Functional Integral Approach
Previously, we primarily examined the relationship between the quantum state of the under-study system and that of the measuring apparatus from the perspective of quantum states, particularly entangled states or relative state. We discussed the evolution equation that governs the evolution of the quantum state space metric of the entangled subsystem relative to another subsystem, as well as its relationship with the Schrödinger equation. Additionally, we explored how to project entangled states onto subspaces using subspace metrics. This approach represents a canonical formulation based on quantum states. In this appendix, we will address this issue from the perspective of relative evolution based on the functional method. The advantage of the functional integral approach is that quantum operators, treated as q-numbers (quantum numbers), can be regarded as ordinary c-numbers (classical numbers) when subjected to mean-field or semi-classical approximation. This facilitates a recovering to the conventional standard quantum theory, making it easier to compare with the standard quantum theories and to discern how this theory surpasses the standard framework.
As previously mentioned, what holds physical meaning is the relationship between the object’s position and the position of the clock pointer. In order to clarify how the position operator $X$ of the object evolves with respect to the position operator $T$ of the clock pointer, we consider the action formulation of this theory
$S_{X}=\int d\tau\left[\frac{1}{2}m_{X}\left(\frac{dX}{d\tau}\right)^{2}-V(X)\right],\quad S_{T}=\frac{1}{2}m_{T}\int d\tau\left[\left(\frac{dT}{d\tau}\right)^{2}\right]$ (61)
Unlike the relative state approach, which does not require any external absolute parameters, the functional method now necessitates a global parameter $\tau$, but the physical results do not depend on it. Both the object’s position $X$ and the clock $T$ evolve relative to this global parameter $\tau$, and in this way, $X$ evolves relative to $T$. In this model, the global parameter $\tau$ can be interpreted as the “proper time” of the laboratory as discussed in section II.
The action for the object is given by the kinetic energy $\frac{1}{2}\left(\frac{dX}{d\tau}\right)^{2}$ minus certain potential energy $V(X)$ with respect to the laboratory proper time $\tau$. We assume that the clock pointer moves uniformly with respect to the laboratory time $\tau$, so the action for the clock consists solely of kinetic energy. For the two mutually entangled systems $|X,T\rangle$ has no interaction with each other, the actions for the object $X$ and the clock $T$ are the direct sums of the actions of two non-interacting systems,
$S[X,T]=S_{X}+S_{T}=\int d\tau\left[\frac{1}{2}m_{X}\left(\frac{dX}{d\tau}\right)^{2}-V(X)+\frac{1}{2}m_{T}\left(\frac{dT}{d\tau}\right)^{2}\right]$ (62)
We observe that, since the total action $S[X,T]$ does not contain a linear term in $T$, the total average energy of the system w.r.t. the internal clock time $T$ is zero
$\frac{\delta S}{\delta T}=\langle E\rangle=0$ (63)
Thus, the total Hamiltonian $H_{X}+H_{T}$ of the entire system w.r.t. the internal clock time equals zero, and the Schrödinger equation governing the evolution of the system’s quantum state $|X,T\rangle$ with respect to $T$ is, strictly speaking, a Wheeler-DeWitt equation
$\left(H_{X}+H_{T}\right)|X,T\rangle=0$ (64)
This equation corresponds to the fact that the entangled state does not evolve w.r.t. $T$ and remain fixed after calibration. Instead, the evolution of the object’s position $X$ relative to the quantum clock $T$ arises because both $X$ and $T$ both evolve with respect to the common laboratory proper time $\tau$. Therefore, to obtain the evolution of $X$ relative to the clock $T$, we only need to perform a semiclassical coordinate transformation $\tau\rightarrow T$ (“semiclassical” means that the operator $T$ is treated using the mean-field/semiclassical approximation, handled as a c-number rather than a q-number and ignore its fluctuation), and then the action can be rewritten as
$S[X,T]=\int dT\left\|\frac{d\tau}{dT}\right\|\left\{\left(\frac{dT}{d\tau}\right)^{2}\left[\frac{1}{2}m_{X}\left(\frac{\delta X}{\delta T}\right)^{2}+\frac{1}{2}m_{T}\right]-V(X)\right\}$ (65)
in which $\left\|\frac{d\tau}{dT}\right\|$ represents the lapse function of time, which is essentially the Jacobian of the coordinate transformation $\tau\rightarrow T$. This transformation formally converts the original integral with respect to $\tau$ into an integral with respect to $T$. Quantizing this theory is equivalent to calculating the partition function
$Z=\int\mathscr{D}X\mathscr{D}Te^{iS[X,T]}$ (66)
By employing the semiclassical approximation, we only consider the semi-classical (S.C.) value of the clock $\langle T\rangle$ and neglect its fluctuation or variance $\langle\delta T^{2}\rangle=\langle T^{2}\rangle-\langle T\rangle^{2}$. In this case, the contribution from the classical path $\langle T\rangle$ of $T$ dominates the path integral over $T$. By taking its average value in the action, the partition function transforms into its semiclassical result
$Z=\int\mathscr{D}X\mathscr{D}Te^{iS[X,T]}\overset{S.C.}{\approx}\int\mathscr{D}Xe^{iS[X\left(\langle T\rangle\right)]}$ (67)
Then, up to an extra constant, we obtain an action that is highly analogous to the action $S_{X}\left[X(\tau)\right]$, with the sole difference being that the laboratory proper time parameter $\tau$ is replaced by the clock time $\langle T\rangle$, and the derivative of $X$ with respect to $\tau$ in the kinetic term is now replaced by a functional derivative of $X$ with respect to $T$
$S[X\left(\langle T\rangle\right)]=\int dT\left[\frac{1}{2}M_{X}\left(\frac{\delta X}{\delta T}\right)^{2}-V(X)+\textrm{const}\right]$ (68)
in which the effective mass $M_{X}$ of the object with respect to the clock time $T$ becomes
$M_{X}=m_{X}\left\|\frac{d\tau}{d\langle T\rangle}\right\|\left(\frac{d\langle T\rangle}{d\tau}\right)^{2}=m_{X}\frac{d\langle T\rangle}{d\tau}$ (69)
Only when the clock runs at the same rate as the laboratory proper time, that is, when the lapse function $\frac{d\langle T\rangle}{d\tau}=1$, does $M_{X}$ equal $m_{X}$. In this sense, mass is not actually an invariant under general coordinate transformations, but merely an invariant under transformations between inertial coordinate systems.
When the quantum fluctuations of the clock can be neglected (under the semiclassical approximation of the clock), the position of the clock’s pointer is represented by a c-number rather than an operator. In this case, the Schrödinger equation governing evolution relative to this clock can be derived as a semiclassical approximation to the Wheeler-DeWitt equation. Therefore, the total Hamiltonian recovers
$H[X\left(\langle T\rangle\right)]\overset{S.C.}{\approx}\frac{1}{2}M_{X}\left(\frac{dX}{dT}\right)^{2}+V(X)+\textrm{const}^{\prime}$ (70)
The Hamiltonian $H_{X}+H_{T}$ in the Wheeler-DeWitt equation now degenerates into the Hamiltonian $H[X\left(\langle T\rangle\right)]$ of the Schrödinger equation with $\langle T\rangle$ as a parameter time. Meanwhile, the solution to the Wheeler-DeWitt equation, the entangled state $|X,T\rangle=\sum_{\tau}C_{\tau}|X(\tau)\rangle\otimes|T(\tau)\rangle$ reduces to the solution of the Schrödinger equation, namely the standard absolute quantum state $|X\left(\langle T\rangle\right)\rangle$. Under the semiclassical approximation, approximating the coordinate operator $T$ of the quantum clock as the parameter $\langle T\rangle$, the quantum state $|X\left(\langle T\rangle\right)\rangle$ is equivalent to projecting the entangled state $|X,T\rangle$ by using the condition projector $\langle\tau^{\prime}|T\rangle=\sum_{\tau}A_{\tau}\delta_{\tau\tau^{\prime}}$, which gives $\langle T|X,T\rangle$ (8), and $\delta_{\tau\tau^{\prime}}=\langle\tau^{\prime}|T(\tau)\rangle$ represents an infinitely precise, zero-width quantum clock eigenstate $|T(\tau)\rangle$ under the semiclassical approximation. This yields $|X\left(\langle T\rangle\right)\rangle=\sum_{\tau}\frac{C_{\tau}}{A_{\tau}}|X(\tau)\rangle$, when the measuring instrument is in the reference state $|T\rangle=\sum_{\tau}A_{\tau}|T(\tau)\rangle$, and hence the relative amplitude of the under-study system being in the state $|X(\tau)\rangle$ is $\frac{C_{\tau}}{A_{\tau}}$.
Compared with the relative state approach, one of the advantages of the functional integral approach is that the effects of the second-order moment quantum fluctuation of the quantum clock $\langle\delta T^{2}\rangle$ (and that of the quantum reference frame fields) is more directly to calculate in its perturbation expansion, which begins to have non-trivial quantum effect in quantum gravity. Once the quantum fluctuations of the clock, $\langle\delta T^{2}\rangle$, cannot be neglected—that is, when the eigenstate of the clock, $\langle\tau^{\prime}|T(\tau)\rangle$, can no longer be approximated semiclassically using a delta function and instead requires consideration of its second-moment quantum fluctuation, such as in the case of a coherent state with finite width—this leads to numerous significant consequences. Firstly, although the average energy of the system, $\langle E\rangle=0$, the fluctuations of the clock, $\langle\delta T^{2}\rangle$, will result in vacuum energy fluctuations, $\langle\delta E^{2}\rangle\neq 0$, providing the correct contribution of vacuum energy to the cosmological constant (Luo2014The, ; Luo2015Dark, ; Luo:2015pca, ; Luo:2019iby, ; Luo:2021zpi, ). Secondly, the Schrödinger equation no longer strictly holds, which causes the constant term in the semiclassical Hamiltonian (70) to cease being a constant. This naturally introduces fluctuation effects akin to some “inertial forces”, as demonstrated in (Anandan:1990fq, ; 2024arXiv240809630L, ).
When this model of a quantum clock with a single degree of freedom is generalized to a quantum spacetime reference frame with multiple degrees of freedom, the aforementioned results are extended in a similar manner: the action $S_{T}$ of the single-degree-of-freedom quantum clock is generalized to a non-linear sigma model with 4-degrees of freedom $X,Y,Z,T$, serving as the quantum spacetime reference frame fields. Meanwhile, the action $S_{X}$ of the object’s position can be generalized to the action of a general quantum field, e.g. $\phi$, representing the quantum system under-study. The global parameter $\tau$ is extended from a single parameter to the four Minkowski coordinate parameters $x,y,z,t$ of the laboratory spacetime, which act as the base space of the non-linear sigma model, while the target space corresponds to the actual spacetime $X,Y,Z,T$. The resulting model, under the semiclassical approximation, yields the effective action of the quantum field $\phi$ on a general curved spacetime background. Furthermore, when considering the effects of second-order quantum fluctuations of the quantum spacetime reference frame fields beyond the semiclassical approximation, the effects of “inertial forces”—namely, gravity—automatically emerge, and, under certain approximation, it takes on the form of Einstein gravity (Luo:2021zpi, ; 2024arXiv240809630L, ; Luo:2025cer, ).